### v2

# PPR: Enhancing Dodging Attacks while Maintaining Impersonation Attacks on Face Recognition Systems

Abstract
Adversarial Attacks on Face Recognition (FR) encompass two types: impersonation attacks and evasion attacks. We observe that achieving a successful impersonation attack on FR does not necessarily ensure a successful dodging attack on FR in the black-box setting. Introducing a novel attack method named Pre-training Pruning Restoration Attack (PPR), we aim to enhance the performance of dodging attacks whilst avoiding the degradation of impersonation attacks. Our method employs adversarial example pruning, enabling a portion of adversarial perturbations to be set to zero, while tending to maintain the attack performance. By utilizing adversarial example pruning, we can prune the pre-trained adversarial examples and selectively free up certain adversarial perturbations. Thereafter, we embed adversarial perturbations in the pruned area, which enhances the dodging performance of the adversarial face examples. The effectiveness of our proposed attack method is demonstrated through our experimental results, showcasing its superior performance.
1 Introduction
Figure 1: An example of comparison between our proposed attack (blue) and the traditional impersonation attack (red) in the black-box setting. The letters ‘D’ and ‘S’ signify that the FR model classifies the two face images as belonging to different identities and belonging to the same identity, respectively.
Thanks to the ceaseless advancements in deep learning, FR has made remarkable strides in achieving exceptional levels of performance(Schroff et al., 2015)(Wang et al., 2018)(Deng et al., 2022). However, the vulnerability of existing FR models to adversarial attacks poses a significant threat to their security.
Hence, there is an urgent need to enhance the performance of adversarial face examples to expose more blind spots in FR models. Consequently, several research endeavors have been directed towards this realm.
A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness(Qiu et al., 2020)(Yang et al., 2021)(Cherepanova et al., 2021)(Hu et al., 2022)(Shamshad et al., 2023), transferability(Zhong & Deng, 2021)(Zhou et al., 2023b)(Zhou et al., 2023a), and physical attack capability(Yin et al., 2021)(Yang et al., 2023)(Li et al., 2023b). These efforts contribute to enhancing the effectiveness of existing attacks using face adversarial examples. However, these studies primarily concentrate on bolstering either impersonation attacks or dodging attacks, overlooking the exploration of the effectiveness of dodging attacks when crafting adversarial face examples using the impersonation attacks.
Unlike image classification, FR is an open-set task. Predicting the class probability of identities in real-world deployment is an extremely challenging task. In practical FR applications, we extract embeddings from two face images using the FR model. Subsequently, the distance between the two embeddings is used to determine whether the images belong to the same identity. If the distance falls below a predefined threshold, the two images are recognized as belonging to the same identity; otherwise, they are classified as different identities. Based the measurement of FR, there exists the samples that can be classified as two different identities in theory. We denote these samples as multi-identity samples. Our experiments have verified the existence of multi-identity samples on common-used face dataset (refer to Appendix A).
In impersonation attacks, the objective is to deceive the FR model into recognizing adversarial face examples as a predefined identity, which requires that the distance between the embeddings of the adversarial face example and the victim image be lower than the threshold. On the other hand, in dodging attacks, the aim is to trick the FR model into recognizing the adversarial face example as identities other than the attacker, which means that the distance between the embeddings of the adversarial face example and the attacker image should be greater than the threshold. The existence of multi-identity samples implies that a successful impersonation attack on FR does not necessarily guarantee a successful dodging attack on FR in theory. Our experiment confirms that the majority of adversarial face examples crafted by traditional impersonation attacks, which can successfully perform impersonation attacks, fail to perform dodging attacks in the black-box setting (see Appendix B).
In practical application scenarios, criminals tend to generate adversarial examples with their own faces to deceive FR models into recognizing them as victims for impersonation attacks, while avoiding being recognized as criminals themselves for dodging attacks to escape legal prosecution. This necessitates the generation of adversarial examples that can perform both impersonation and dodging attacks.
To improve the dodging performance of the adversarial face examples without degrading the impersonation performance, we propose a novel attack method, termed as Pre-training Pruning Restoration Attack (PPR). Specifically, inspired by the pruning methods in model compression, we prune the pre-trained adversarial face examples to sparsify the adversarial face examples. Subsequently, we add adversarial perturbations that favor dodging attacks on the pruned regions to improve the dodging performance of the adversarial face examples (i.e. the operation of Restoration). After pruning and restoration, the impersonation performance of the adversarial face examples can be maintained and the dodging performance of the adversarial face examples can be improved.
Figure 1 illustrates a comparison between our proposed attack and the traditional impersonation attack. The adversarial face example generated by our proposed attack method successfully evades identification as the attacker in contrast with the adversarial face example crafted by the traditional impersonation attack method is identified as the attacker. The adversarial face example generated by our proposed attack method successfully evades identification as the attacker, while the adversarial face example crafted by the traditional impersonation attack method is identified as the attacker. Note that both the adversarial face examples crafted by our proposed attack method and the traditional impersonation attack method are capable of being identified as the victim image.
Our main contributions are summarized as follows:
  • We verify the universality of multi-identity samples among adversarial face examples crafted by traditional impersonation attacks and demonstrate that a successful impersonation attack does not necessarily equate to a successful dodging attack on FR systems.
  • We propose a novel adversarial attack method called PPR based on adversarial example pruning. By pruning the adversarial face examples and subsequently introducing adversarial perturbations favoring dodging attacks, we effectively enhance the dodging performance. The approach allows us to maintain the impersonation performance while significantly improving the dodging performance of the adversarial face examples.
  • Extensive experiments demonstrate that our proposed method achieves superior performance compared to the baseline adversarial attack methods.
2 Related Works
2.1 Adversarial Attacks
The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images for the purpose of deceiving machine learning systems and causing them to make mistakes(Szegedy et al., 2014)(Goodfellow et al., 2015). The existence of adversarial examples poses a significant threat to the security of current machine learning systems. Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems(Long et al., 2022)(Zhang et al., 2022)(Liang et al., 2023)(Wei et al., 2023)(Lu et al., 2023).
To improve the performance of black-box adversarial attacks, DI is proposed by (Xie et al., 2019) to randomly transform adversarial examples in each iteration. VMI-FGSM, proposed by (Wang & He, 2021), employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance. SSA, as proposed by (Long et al., 2022), transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them. (Wang et al., 2023b) propose SIA. SIA applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculations. (Wang et al., 2023a) propose BSR. BSR divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation.
2.2 Face Recognition
FR is an open-set task, meaning we cannot predict the class of a user upon deploying an FR model. As a result, relying solely on the maximum probability value of the output of the neural network to determine the class the of a user becomes impractical. Instead, when deploying the FR model, we extract the embeddings of two images and compute the similarity between them. If the similarity exceeds a predefined threshold, we conclude that the two images belong to the same identity; otherwise, they belong to different identities.
Recent research on FR primarily focuses on reducing the intra-class distance and increasing the inter-class distance of embeddings extracted by FR models. (Schroff et al., 2015) introduce FaceNet that employs triplet loss to map face images into a space where the distance represents the dissimilarity between them. (Wang et al., 2018) develop CosFace, which utilizes the large-margin cosine loss to enhance FR model performance. Additionally, (Deng et al., 2022) introduce ArcFace that incorporates the additive angular margin loss to train FR models that extract more discriminative features.
2.3 Adversarial Attacks on Face Recognition
Adversarial attacks on FR can be mainly classified into two categories: restricted attacks and unrestricted attacks.
Restricted attacks on FR are the attacks that generate adversarial examples in a restricted bound (e.g. $L_{p}$ bound). Because FR belongs to the open-set task, the embedding-level loss is more suitable. Therefore, (Zhong & Deng, 2019) proposed FIM that use embedding-level loss to craft the adversarial face examples. (Zhong & Deng, 2021) proposed DFANet that performs dropout on the feature maps of the convolutional layers to improve the transferability. (Yang et al., 2021) propose TIP-IM to improve the face encryption. TIP-IM use MMD(Borgwardt et al., 2006) loss to improve the quality of encrypted face images and use the greedy insertion to select the optimal victim images. (Zhou et al., 2023b) propose BPFA to improve the transferability of adversarial attacks on FR by adding beneficial perturbations(Wen & Itti, ) on the features maps of the FR models to obtain effect that is similar to hard samples augmentation.
The unrestricted adversarial attacks on FR are the attacks that generate adversarial examples without the restriction of a predefined bound. The unrestricted adversarial attacks on FR mainly focus on physical attacks(Xiao et al., 2021)(Yang et al., 2023)(Li et al., 2023a), attribute editing(Jia et al., 2022) and generating adversarial examples based on makeup transfer(Yin et al., 2021)(Hu et al., 2022)(Shamshad et al., 2023).
3 Methodology
3.1 Problem Formulation
In the context of our study, $\mathcal{F}^{vct}(x)$ denotes the FR model used by the victim to extract the embedding from a face image $x$. We refer to $x^{s}$ and $x^{t}$ as the attacker and victim images, respectively. The objective of the impersonation attacks explored in our research is to manipulate $\mathcal{F}^{vct}$ into misclassifying $x^{adv}$ as $x^{t}$, while ensuring that $x^{adv}$ bears a close visual resemblance to $x^{s}$. To be more specific, the objective can be expressed as follows:
$\begin{gathered}x^{adv}=\mathop{\arg\min}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (1)
where $\mathcal{D}$ refers to a predefined distance metric, while $\epsilon$ specifies the maximum magnitude of permissible perturbation. By contrast, the objective of the dodging attacks proposed in this study is to render $\mathcal{F}^{vct}(x)$ unable to identify $x^{adv}$ as $x^{s}$, while simultaneously ensuring that $x^{adv}$ bears a visual resemblance to $x^{s}$. In a manner similar to impersonation attacks, the objective of dodging attacks can be formulated as follows:
$\begin{gathered}x^{adv}=\mathop{\arg\max}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (2)
Due to the differing objectives of the impersonation and dodging attacks, the calculation methods for the ASR also vary between the two attack types. When it comes to impersonation attacks on FR, the ASR can be computed as:
${\rm ASR}^{i}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)<t^{i}\right)}{N_{p}}$ (3)
where $N_{p}$ refers to the total number of face pairs and $t^{i}$ represents the impersonation attack threshold. For dodging attacks on FR systems, the ASR can be computed via the following formula:
${\rm ASR}^{d}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)>t^{d}\right)}{N_{p}}$ (4)
where the value of $t^{d}$ represents the threshold for dodging attacks.
3.2 Exploring the Impersonation and Dodging Attack on Face Recognition
In most cases, the victim model $\mathcal{F}^{vct}$ is not accessible to the attacker, making it impossible to optimize objectives as stated in Equation 1 and Equation 2 directly. To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversaries that can be transferred to the victim model for an effective attack(Yuan et al., 2022)(Zhang et al., 2022)(Naseer et al., 2023)(Wang et al., 2023b)(Wang et al., 2023a).
For impersonation attacks, the loss can be formulated as follows:
$\mathcal{L}^{i}=\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{t}\right)\right)\|^{2}_{2}$ (5)
where $\phi(x)$ represents the operation that normalizes $x$. $x^{adv}$ is the adversarial example which is initialized with the same value as $x^{s}$. Dodging attacks can be formulated with the following loss function:
$\mathcal{L}^{d}=-\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{s}\right)\right)\|^{2}_{2}$ (6)
As the FR task is an open-set task, it is not possible to predict the classes of users during the practical deployment of the FR model. Therefore, we need to compare the similarity between two face images in order to discern whether they depict the same identity or not. Based on the identification method in face recognition, multi-identity samples exist theoretically. Our experiments verify the existence of such samples among benign face images. The existence of multi-identity samples raises a question:
To this end, we generated face adversarial samples using the traditional impersonation attack and evaluate its dodging ASR (refer to Appendix B). Our experiments demonstrate that the majority of adversarial face examples crafted by traditional impersonation attack cannot achieve a successful dodging attack in the black-box setting.
Nonetheless, in real-world adversarial attacks, attackers do not want the adversarial face examples to be recognized as themselves, as this could lead to legal consequences. Hence, it is crucial to research attack techniques that can execute both impersonation and dodging attacks simultaneously. Traditional impersonation attacks on FR systems have shown a remarkably high level of impersonation ASR in black-box settings. Therefore, our objective is to enhance the dodging performance while maintaining the impersonation effectiveness of traditional impersonation attacks.
To accomplish this objective, a straightforward approach is to generate adversarial face examples by employing a Lagrangian attack strategy that utilizes the following loss function:
$\mathcal{L}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$ (7)
However, due to the conflict between the optimization between $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, there exists a trade-off between the performance of impersonation and dodging performance leading to subpar performance (See Section 4.2). If we alliviete the trade-off, we can achieve a better dodging performance while maintaining the impersonation performance.
3.3 Pre-training Pruning Restoration Attack
To accomplish this objective, a straightforward approach is to fine-tune the adversarial face examples generated by the Lagrangian attack with a lower $\lambda$ value in order to enhance the performance of dodging attacks. However, this method does not enhance the dodging attack performance without compromising the impersonation attack performance (see Fine-tuning in Table 4). We contend that this issue arises because the newly introduced adversarial perturbation that favors dodging attacks ends up disrupting the pre-existing adversarial perturbation patterns. While it may improve the performance of dodging attacks, it inevitably diminishes the performance of impersonation attacks. To address this, we introduce new adversarial perturbations favoring dodging attacks in regions where the original perturbations have not been added. However, identifying suitable areas for these new perturbations is challenging due to their scarcity. Therefore, drawing inspiration from prior works in model pruning(Ma et al., 2023)(Frantar & Alistarh, 2023), we propose incorporating pruning into the realm of adversarial face examples. The aim is to free up some model parameters while minimizing the decrease in the performance of impersonation attacks, thereby creating space to introduce perturbations that facilitate dodging attacks.
In order to prune the adversarial perturbation, our initial step is to assess its importance by identifying and sparsifying the less important regions. To estimate this importance, we propose utilizing the magnitude of the adversarial perturbation as a measure. A lower magnitude implies a lesser impact on the performance of the adversarial examples generated after sparsification.
Let the adversarial examples be $x^{adv}$. The formula to calculate the importance can be express as:
$\mathcal{I}=|x^{adv}-x^{s}|$ (8)
where $\mathcal{I}\in\mathbb{R}^{\rm CHW}$. ${\rm C}$, ${\rm H}$, and ${\rm W}$ are the channel number, hight, width of the face images, respectively.
Let $\kappa$ be the sparsity ratio for pruning the adversarial face examples that measures the ratio of adversarial perturbations to be set into zero. Let $s$ be the number of adversarial perturbation elements:
$s={\rm CHW}$ (9)
We arrange the elements in flattened vector of $\mathcal{I}$ in ascending order (from the lowest to the highest):
$\mathcal{Q}={\rm Sort}\left(\Psi\left(\mathcal{I}\right)\right)$ (10)
where $\Psi$ is the flatten operation.
Let $\mathcal{W}$ be the set of the elements of the adversarial perturbations to be pruned. Given the importance calculation method for pruning, the value of $\mathcal{W}$ can be calculated as follows:
$\mathcal{W}=\mathcal{Q}\left[:\kappa s\right]$ (11)
where the colon denotes the slice operation to obtain the first $\kappa s$ elements.
The mask for pruning can be obtained using $\mathcal{W}$:
$\mathcal{M}_{i,j,k}=\begin{cases}0,&\text{if }\mathcal{I}_{i,j,k}\in\mathcal{W}\\ 1,&\text{if }\mathcal{I}_{i,j,k}\notin\mathcal{W}\end{cases}$ (12)
By utilizing the mask, we can apply the following formula to prune the adversarial example:
$\bar{x}^{adv}=x^{s}+\left(x^{adv}-x^{s}\right)\odot\mathcal{M}$ (13)
Once the adversarial face examples are pruned, we can proceed to restore the adversarial perturbations in the pruned regions. The mask representing the regions for restoring the adversarial examples can be denoted as:
$\mathcal{A}=1-\mathcal{M}$ (14)
For both the pre-training and restoration stages, various adversarial attacks can be employed. In the following, we will illustrate the pre-training and restoration stages of of our proposed method by using the Lagrangian attack as an example. In the stage of pre-training, we utilize Equation 7 as the pre-training loss $\mathcal{L}^{p}$ to craft the adversarial face examples:
$x^{adv}_{t}=x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)$ (15)
where $t$ is the iteration of optimization process of adversarial examples, $\beta$ is the step size when optimizing the adversarial face examples in the pre-training stage.
Once the adversarial examples are pre-trained, we utilize our proposed method of pruning adversarial face examples to free up some adversarial perturbations. During the restoration stage, we restore the adversarial face examples in the previously pruned region using the following loss function:
$\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$ (16)
where $\tilde{\lambda}$ is a weight which is lower than $\lambda$ that is objective for crafting adversarial face examples that favor dodging attacks.
$x^{adv}_{t}=x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)$ (17)
where $\gamma$ is the step size when optimizing the adversarial face examples in the restoration stage. The algorithm of our proposed method based on Lagrangian attack are illustrated in Algorithm 1.
Algorithm 1 Pre-training Pruning Restoration Attack Based on Lagrangian
0:  Negative face image pair $\{x^{s},x^{t}\}$, the step size of the adversarial perturbations $\beta$, the number of the iterations of pre-training stage $n$, the maximum number of iterations $m$, maximum allowable perturbation magnitude $\epsilon$, the surrogate FR model $\mathcal{F}$.
0:  An adversarial face example $x^{adv}_{m}$
1:  $x^{adv}_{0}=x^{s}$
2:  for $t=1,...,n$ do
3:     $\mathcal{L}^{p}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Pre-training Stage
4:     $x^{adv}_{t}=x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)$
5:     $x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{t}\right)$
6:  end for
7:  $\mathcal{I}=|x^{adv}_{n}-x^{s}|$triangleright Pruning Stage
8:  $\mathcal{Q}={\rm Sort}\left(\Psi\left(\mathcal{I}\right)\right)$
9:  $\mathcal{W}=\mathcal{Q}\left[:\kappa s\right]$
10:  Get $\mathcal{M}$ by Equation 12.
11:  $x^{adv}_{n}=x^{s}+\left(x^{adv}_{n}-x^{s}\right)\odot\mathcal{M}$
12:  $\mathcal{A}=1-\mathcal{M}$
13:  for $t=n+1,...,m$ do
14:     $\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Restoration Stage
15:     $x^{adv}_{t}=x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)$
16:     $p=\mathcal{A}\odot\left(x^{adv}_{t}-x^{adv}_{n}\right)$
17:     $x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{n}+p\right)$
18:  end for
4 Experiments
4.1 Experimental Setting
Datasets. We have opted to use the LFW dataset (Huang et al., 2007) and the CelebA-HQ dataset (Karras et al., 2018) for our experiments. LFW serves as an unconstrained face dataset for FR, while CelebA-HQ consists of high-quality images. These two datasets are widely employed in research on adversarial attacks in FR (Xiao et al., 2021; Yin et al., 2021; Hu et al., 2022). The LFW and CelebA-HQ datasets utilized in our experiments are identical to those employed in previous studies (Zhou et al., 2023b, a).
Table 1: Comparisons of dodging ASR results for attacks on the LFW and CelebA-HQ datasets. The surrogate models are presented in the first column, and the victim models are listed in the second row.
LFW CelebA-HQ
Surrogate Model Attack IR152 FaceNet MF IR152 FaceNet MF
IR152 DI 95.4 / 100.0 5.8 / 11.3 0.2 / 0.8 87.9 / 100.0 8.4 / 16.4 0.4 / 1.2
VMI 92.7 / 100.0 17.3 / 32.5 1.2 / 9.5 92.2 / 100.0 14.3 / 27.9 1.2 / 4.1
SSA 78.8 / 100.0 5.7 / 22.5 0.9 / 10.6 83.8 / 99.9 7.2 / 19.6 0.4 / 5.6
DFANet 98.9 / 100.0 1.4 / 4.2 0.0 / 0.3 98.9 / 100.0 2.3 / 6.0 0.0 / 0.4
SIA 81.7 / 100.0 13.0 / 37.5 0.8 / 8.9 78.4 / 100.0 13.2 / 35.6 0.7 / 7.4
BSR 52.4 / 100.0 5.3 / 17.6 0.1 / 1.5 48.5 / 99.9 5.4 / 18.0 0.3 / 1.9
BPFA 92.6 / 100.0 1.7 / 7.3 0.0 / 1.2 90.4 / 100.0 2.1 / 8.1 0.1 / 0.8
FaceNet DI 5.3 / 10.3 99.8 / 99.9 3.1 / 10.3 1.5 / 3.1 99.4 / 99.9 1.8 / 4.7
VMI 9.7 / 14.3 99.8 / 99.9 6.2 / 13.2 3.1 / 7.1 99.3 / 99.8 3.6 / 9.3
SSA 6.0 / 14.0 97.5 / 99.9 6.6 / 26.2 2.0 / 5.5 96.9 / 99.7 4.2 / 14.6
DFANet 1.6 / 3.3 99.8 / 99.9 0.4 / 2.7 0.5 / 2.6 99.1 / 100.0 0.8 / 4.1
SIA 11.2 / 20.6 99.5 / 99.9 8.7 / 21.2 4.0 / 8.9 99.4 / 99.9 5.4 / 13.7
BSR 12.2 / 19.2 98.6 / 99.9 9.0 / 17.8 4.6 / 10.1 98.8 / 99.9 5.3 / 14.1
BPFA 4.7 / 16.8 98.6 / 100.0 1.6 / 15.0 1.1 / 4.2 99.0 / 100.0 0.6 / 5.1
MF DI 2.2 / 7.3 18.2 / 36.4 99.2 / 100.0 0.1 / 2.5 12.1 / 31.3 95.2 / 100.0
VMI 1.0 / 2.8 8.4 / 20.9 99.7 / 100.0 0.2 / 0.4 5.2 / 15.0 98.2 / 100.0
SSA 0.7 / 4.1 6.1 / 23.5 98.3 / 100.0 0.0 / 0.6 3.9 / 18.5 93.3 / 100.0
DFANet 0.2 / 1.0 1.5 / 5.8 99.6 / 100.0 0.0 / 0.2 1.1 / 7.7 99.1 / 100.0
SIA 1.0 / 5.9 10.6 / 36.6 98.4 / 100.0 0.1 / 2.4 9.0 / 24.4 96.3 / 100.0
BSR 0.4 / 1.5 3.7 / 14.7 84.9 / 100.0 0.1 / 0.6 2.9 / 12.6 77.6 / 100.0
BPFA 0.9 / 4.1 4.6 / 20.4 97.7 / 100.0 0.0 / 2.3 4.0 / 20.4 96.2 / 100.0
[…]
For calculating the attack success rate in impersonation and dodging attacks, we choose the thresholds based on FAR@0.001 on the entire LFW dataset.
Attack Setting. Without any particular emphasis, we set the maximum allowable perturbation magnitude, denoted as $\epsilon$, to 10. This value is based on the $L_{\infty}$ norm bound, considering pixel values within the range of $[0,255]$. Furthermore, we specify the maximum number of iterative steps as 200. Moreover, both $\beta$ and $\gamma$ are set to 1.0.
Evaluation Metrics. We employ the attack success rate (ASR) to assess the performance of various attacks. ASR represents the ratio of successfully attacked adversarial face examples.
Baseline methods Our proposed attack is a restricted attack method that aims to enhance the robustness of FR models. It is not rational to compare our proposed method with the unrestricted attacks which do not limit the magnitude of the adversarial perturbations. Therefore, we choose restricted attacks on FR that aim to enhance the robustness of FR models(Zhong & Deng, 2021)(Zhou et al., 2023b) and transfer attacks(Xie et al., 2019)(Long et al., 2022)(Wang et al., 2023b)(Wang et al., 2023a) as our baseline.
Table 2: Comparisons of dodging ASR with multi-task attacks on the LFW dataset. The models in the first row are the victim models.
[…]
IR152 FaceNet MF
Lagrangian 3.9 26.5 100.0 26.0
Lagrangian + ours 7.3 36.4 100.0 26.6
DA 11.0 35.6 99.1 37.4
DA + ours 17.5 44.9 99.4 37.8
4.2 The Trade-off Between the Impersonation Attacks and Dodging Attacks
Owing to the inherent conflict during the optimization process of impersonation and dodging losses, there exists a trade-off between impersonation and dodging performance. We craft adversarial face examples using Lagrangian attack and our proposed PPR on LFW based on DI. The average black-box results are demonstrated in Figure 2.
Figure 2: The trade-off between the performance of impersonation attack and dodging attack of adversarial examples.
Figure 2 illustrates that the inclusion of our proposed method into the Lagrangian attack results in reduced the trade-off. It is not hard to see the reason. The pruning operation of our proposed PPR serves to sparsify the adversarial perturbations while preserving the impersonation performance. On the other hand, the restoration operation tends to introduce adversarial perturbations in the pruned areas, specifically favoring dodging attacks. These operations effectively enhance the dodging attack performance while maintaining the impersonation attack performance, ultimately mitigating the trade-off.
4.3 Comparison Study
We compare our proposed attack method with the baseline attacks on multiple FR models and datasets. For our proposed method, we choose the Lagrangian attack as the attack for both the stages of pre-training and restoration. The results are shown in Table 1. The numbers before the slash represent the results of the baseline attack, which refers to the attack method mentioned in the seconds column that solely utilizes the traditional impersonation attacks to create the adversarial examples. On the other hand, the numbers after the slash correspond to the results obtained by combining the baseline attack with our proposed PPR.
Table 1 shows that after being added our proposed attack method, the dodging ASR of the adversarial attacks improve a lot. Note that the average black-box impersonation ASRs of the baseline attacks in Table 1 increase after incorporating our proposed attack method. This demonstrate the effectiveness of our proposed method in improving the dodging attack performance while maintain the impersonation attack performance.
[…]
In practical application scenarios, victims can employ adversarial robust models to defend against adversarial attacks. Consequently, it becomes crucial to evaluate the performance of adversarial attacks on these robust models. In this study, we generate adversarial examples on the LFW dataset using MF as the surrogate model and assess the performance of various attacks on the adversarial robust models. The results are presented in Table 3.
Table 3 illustrates that the inclusion of our proposed method leads to improvements in both dodging and impersonation performance. These results serve as evidence of the effectiveness of our proposed method on adversarial robust models.
Table 3: Comparisons of dodging ASR on LFW with adversarial robust models as victim models.
[…]
DI 2.3 7.8 0.5 19.8
DI + ours 7.2 17.1 4.7 20.3
VMI 0.8 4.5 0.3 15.8
VMI + ours 3.3 8.9 3.0 16.2
SSA 0.6 2.9 0.1 14.7
SSA + ours 3.9 9.7 3.0 14.8
DFANet 0.2 0.8 0 8.1
DFANet + ours 0.3 1.3 0.1 8.3
SIA 1.4 6 0.2 17.5
SIA + ours 5.6 15.5 5.6 17.6
BSR 0.4 1.9 0 6.4
BSR + ours 0.7 3.5 0.4 6.6
BPFA 0.3 1.8 0.1 13.1
BPFA + ours 4.1 8.4 4.2 13.4
4.4 Ablation Study
To delve deeper into the properties of our proposed attack method, we conducted an ablation experiment using DI as the Baseline attack, with MF serving as the surrogate model on the LFW dataset. To confirm the effectiveness of our pruning method, we employed the Random Zeroing (RZ) method, which randomly sets adversarial perturbations to zero. We applied this method and our pruning method to free up 20% of the adversarial perturbations. For Fine-tuning, we employed the Lagrangian attack as the method to further optimize the adversarial examples crafted by the Lagrangian attack. The dodging attack ASR and average black-box impersonation ASR results are shown in Table 4.
Table 4: Comparisons of ASR results of dodging attack and impersonation attacks on the LFW dataset. * denote the method that only use the impersonation loss to craft the adversarial face examples.
[…]
IR152 FaceNet MF
Baseline 2.2 18.2 99.2 25.6
Lagrangian 3.9 26.5 100.0 26.0
Fine-tuning 3.4 23.2 99.9 25.8
RZ 0.6 4.4 94.8 15.1
Pruning 3.4 24.8 100.0 25.4
PPR 6.8 35.5 100.0 25.9
Table 4 demonstrates that our proposed pruning method for adversarial examples achieves a significantly smaller decrease in ASR than RZ after pruning 20% of adversarial perturbations, indicating the effectiveness of our proposed approach. After pruning and restoration, adversarial perturbations using our proposed Pruning method, both impersonation and dodging ASR of the crafted adversarial face examples are recovered and higher than the Baseline attack method. These results demonstrate the effectiveness of our proposed PPR in improving the dodging performance of adversarial attacks on FR without compromising the impersonation attack performance.
Figure 3: The dodging ASR in different sparsity ratio.
4.5 Hyperparameter Sensitivity Study
The sparsity ratio quantifies the proportion of adversarial perturbations that are allowed to be discarded during the pruning stage. This ratio greatly impacts the performance of our proposed attack method. Hence, we conducted a sensitivity study on the sparsity ratio to analyze its effect on performance of the algorithm. We selected the DI-based Lagrangian attack method as the Baseline for pre-training and performed a hyperparameter sensitivity study on LFW with FaceNet as the surrogate model. We adjusted the value of $\tilde{\lambda}$ to ensure that the average black-box impersonation ASR results were within a 0.8% absolute difference compared to the corresponding Baseline. The dodging ASR results are illustrated in Figure 3.
Figure 3 illustrates that the dodging ASR of our proposed method initially increases and then decreases as the sparsity ratio increases. It is not hard to see the reason; When the sparsity ratio increases, a greater number of adversarial perturbations are pruned, creating more empty regions for the adversarial perturbations that favor dodging attacks in the restoration stage. If the sparsity ratio is set to a too high value, an excessive number of adversarial perturbations are allowed to be freed up, resulting in a degradation of performance for the pre-trained adversarial examples. Consequently, the performance of adversarial face examples will decrease.
5 Conclusion
In this paper, we delve into the issue of multi-identity samples present among adversarial face examples. Our research reveals that traditional impersonation attacks often involve multi-identity samples and the success of an traditional impersonation attack may not necessarily imply success in dodging attacks in the black-box setting. In order to improve dodging performance without compromising impersonation performance, we proposed a novel attack known as PPR. PPR involves pruning the adversarial examples to release some of the adversarial perturbations while tending to maintain the performance of the adversarial examples. We then restore the adversarial perturbation in the pruned area, which favors dodging attack, to enhance the performance of dodging attack without harming the performance of the impersonation attack. Our extensive experiments demonstrate the effectiveness of our proposed attack method.
6 Potential Broader Impact
This paper showcases research dedicated to advancing the field of Machine Learning. The proposed attack method could jeopardize the security of FR systems. Our aim is to raise awareness through our proposed attack method and reinforce the robustness of FR systems.
Appendix A Multi-identity Samples among the Benign Face Images
Multi-identity samples are intriguing samples that can be classifies as multiple classes in FR. In this subsection, we will explore the existence of multi-identity samples among the benign face images (i.e. the face images without being added adversarial perturbations).
We randomly select negative face pairs from the entire LFW dataset. Subsequently, we use MF as our FR model to extract the embeddings of the face images in each face pair and calculate the cosine similarity between the two images. If the cosine similarity surpasses the threshold, both images in the pair are classified as belonging to multiple identities, indicating that they are multi-identity samples. Our findings demonstrate the presence of multi-identity samples among the benign face images, as illustrated in Figure 4.
Figure 4 illustrates that the multi-identity samples among the benign face images closely resemble the appearances of the identities they are classified into, suggesting they are hard samples(Zhou et al., 2023b)(Huang et al., 2020).
To analyze the cause of this phenomenon, we need to consider the properties of both the multi-identity samples and the FR model. Commonly-used FR models are well-trained and capable of correctly classifying the majority of benign face images. However, there are some benign face images that the FR model fails to classify accurately, and these samples are referred to as hard samples. Multi-identity samples are a specific type of hard sample known as hard negative samples. Typically, hard negative samples exhibit a similar appearance(Zeng et al., 2020), indicating that the multi-identity samples among the benign face images share a resemblance.
Figure 4: The Illustration of the multi-identity samples. The images in the first row and third row are the images of different identities. The image in the second row are the images of the multi-identities samples that will be recognized as the identities both top and below it by the FR model.
Appendix B Universality of Multi-identity Samples among Adversarial Face Examples
The traditional impersonation attacks craft adversarial face example only use the impersonation loss (e.g. Equation 5). In this subsection, we will investigate the ratio of multi-identity samples and evaluate the effectiveness of traditional impersonation attacks in terms of dodging attacks. We utilize the multi-identity sample ratio (MSR) to gauge the proportion of multi-identity samples that can be recognized as both the attacker and victim identities in the adversarial face examples capable of executing successful impersonation attacks. MSR be calculate as:
${\rm MSR}=\frac{|\mathcal{S}_{1}\cap\mathcal{S}_{2}|}{|\mathcal{S}_{2}|}$ (18)
where $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are the set of the adversarial examples that are identified as the attacker and victim identities, respectively. Therefore, the $\mathcal{S}_{2}$ is the set of the adversarial face examples that can perform successful impersonation attack. Equation 3 can be rewritten as:
${\rm ASR}^{i}=\frac{|\mathcal{S}_{2}|}{N_{p}}$ (19)
We evaluate the MSR, impersonation attack ASR and dodging attack ASR using MF as the surrogate model on the LFW dataset. The results are shown in Table 5.
Table 5: The results (%) on the LFW dataset. The models on the first row are the victim models.
Attack Metric IR152 FaceNet MF
DI MSR 95.6 75.7 0.8
[…]
VMI MSR 97.0 88.6 0.3
[…]
SSA MSR 98.6 92.8 1.5
[…]
DFANet MSR 98.6 96.6 0.4
[…]
SIA MSR 97.4 83.9 1.6
[…]
BSR MSR 98.1 96.8 15.1
[…]
BPFA MSR 96.2 92.1 2.3
[…]
Table 5 demonstrate that most of the crafted adversarial face examples are multi-identity samples when the victim model are the black-box model. This indicates that most adversarial face examples generated through traditional impersonation attacks are unable to attain a successful dodging attack in the black-box setting. Nevertheless, in the white-box setting, the majority of adversarial face examples capable of executing successful impersonation attacks also demonstrate success in dodging attacks.
To analyze the reason, we need to consider the metric used to determine whether two face images belong to the same identity in FR. FR is an open-set task, so we utilize the distance in the embedding space to make this determination. Figure 5 illustrates two decision boundaries for each FR model, one for attacker identity and one for victim identity.
In the white-box setting, if we generate adversarial examples using Equation 5, these examples can enter a space in which they are identified as the victim identity but not as the attacker identity. However, the decision boundary of the black-box model differs from that of the surrogate model. In the black-box setting, most adversarial examples are found between the decision boundaries of the black-box model, resulting in the majority of adversarial examples crafted using Equation 5 being multi-identity samples. This demonstrates that the adversarial face examples are positioned near the decision boundary, aligning with findings from prior research of adversarial attacks on image classification(Cao & Gong, 2017).
Figure 5: Illustration of the decision boundaries of FR models. The star in the left and right are the initial image of the adversarial face examples (i.e. the attacker image) and the final adversarial example, respectively. The solid and dashed lines are the decision boundaries for the surrogate and black-box models, respectively. For both the surrogate and black-box models, the lines in the left and right are the decision boundary of the victim and attacker, respectively.

### v3

@@
-PPR: Enhancing Dodging Attacks while Maintaining Impersonation Attacks on Face Recognition Systems
+Rethinking Impersonation and Dodging Attacks on Face Recognition Systems
 Abstract
-Adversarial Attacks on Face Recognition (FR) encompass two types: impersonation attacks and evasion attacks.
-We observe that achieving a successful impersonation attack on FR does not necessarily ensure a successful dodging attack on FR in the black-box setting.
-Introducing a novel attack method named Pre-training Pruning Restoration Attack (PPR), we aim to enhance the performance of dodging attacks whilst avoiding the degradation of impersonation attacks.
-Our method employs adversarial example pruning, enabling a portion of adversarial perturbations to be set to zero, while tending to maintain the attack performance.
-By utilizing adversarial example pruning, we can prune the pre-trained adversarial examples and selectively free up certain adversarial perturbations.
-Thereafter, we embed adversarial perturbations in the pruned area, which enhances the dodging performance of the adversarial face examples.
-The effectiveness of our proposed attack method is demonstrated through our experimental results, showcasing its superior performance.
+Face Recognition (FR) systems can be easily deceived by adversarial examples that manipulate benign face images through imperceptible perturbations.
+Adversarial attacks on FR encompass two types: impersonation (targeted) attacks and dodging (untargeted) attacks.
+Previous methods often achieve a successful impersonation attack on FR; However, it does not necessarily guarantee a successful dodging attack on FR in the black-box setting.
+In this paper, our key insight is that the generation of adversarial examples should perform both impersonation and dodging attacks simultaneously.
+To this end, we propose a novel attack method termed as Adversarial Pruning (Adv-Pruning), to fine-tune existing adversarial examples to enhance their dodging capabilities while preserving their impersonation capabilities.
+Adv-Pruning consists of Priming, Pruning, and Restoration stages.
+Concretely, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations, identifying and releasing those with minimal impact on absolute model output variances.
+Then, Biased Gradient Adaptation is presented to adapt the adversarial examples to traverse the decision boundaries of both the attacker and victim by adding perturbations favoring dodging attacks on the vacated regions, preserving the prioritized features of the original perturbations while boosting dodging performance.
+As a result, we can maintain the impersonation capabilities of original adversarial examples while effectively enhancing dodging capabilities.
+Comprehensive experiments demonstrate the superiority of our method compared with state-of-the-art adversarial attacks.
 1 Introduction
-Figure 1: An example of comparison between our proposed attack (blue) and the traditional impersonation attack (red) in the black-box setting.
-The letters ‘D’ and ‘S’ signify that the FR model classifies the two face images as belonging to different identities and belonging to the same identity, respectively.
-Thanks to the ceaseless advancements in deep learning, FR has made remarkable strides in achieving exceptional levels of performance(Schroff et al., 2015)(Wang et al., 2018)(Deng et al., 2022).
+Figure 1: Top: previous methods that achieve a successful impersonation attack on FR cannot guarantee a successful dodging attack on FR in the black-box setting.
+In contrast, we present Adv-Pruning, including Priming, Pruning, and Restoration Stages, to perform both impersonation and dodging attacks simultaneously.
+Bottom (left): natural Multi-identity Samples (MS).
+Bottom (right): the dodging Attack Success Rate (%) between the previous methods and Adv-Pruning on multiple models.
+Thanks to the ceaseless advancements in deep learning, Face Recognition (FR) has achieved exceptional performance [41; 49; 9; 1; 2; 26].
 However, the vulnerability of existing FR models to adversarial attacks poses a significant threat to their security.
 Hence, there is an urgent need to enhance the performance of adversarial face examples to expose more blind spots in FR models.
-Consequently, several research endeavors have been directed towards this realm.
-A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness(Qiu et al., 2020)(Yang et al., 2021)(Cherepanova et al., 2021)(Hu et al., 2022)(Shamshad et al., 2023), transferability(Zhong & Deng, 2021)(Zhou et al., 2023b)(Zhou et al., 2023a), and physical attack capability(Yin et al., 2021)(Yang et al., 2023)(Li et al., 2023b).
-These efforts contribute to enhancing the effectiveness of existing attacks using face adversarial examples.
-However, these studies primarily concentrate on bolstering either impersonation attacks or dodging attacks, overlooking the exploration of the effectiveness of dodging attacks when crafting adversarial face examples using the impersonation attacks.
-Unlike image classification, FR is an open-set task.
-Predicting the class probability of identities in real-world deployment is an extremely challenging task.
-In practical FR applications, we extract embeddings from two face images using the FR model.
+As a result, several research endeavors have been directed towards this realm.
+A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness [40; 62; 7; 19; 42], transferability [70; 29; 72; 71], and physical attack capability [64; 63; 28].
+These efforts contribute to enhancing the effectiveness of adversarial attacks on FR.
+Nevertheless, these studies primarily concentrate on bolstering either impersonation attacks or dodging attacks, overlooking the exploration of the effectiveness of dodging attacks when crafting adversarial face examples using impersonation attacks.
+In real-world deployment contexts, individuals with malicious intent are prone to creating adversarial face examples incorporating their own facial features to manipulate FR systems to mistakenly identify them as pre-defined victims during impersonation attacks.
+Concurrently, the individuals strive to evade accurate identification as perpetrators, thereby circumventing detection and preventing legal accountability.
+This requires the creation of adversarial examples capable of executing both impersonation and dodging attacks simultaneously.
+In the realm of adversarial attacks on image classification, a successful impersonation attack typically implies a successful dodging attack.
+However, FR is an open-set task [49; 9], which is quite different from image classification.
+In the real-world deployment of FR systems, accurately predicting the class probability of identities presents an extreme challenge.
+Therefore, we extract embeddings from two face images using the FR model.
 Subsequently, the distance between the two embeddings is used to determine whether the images belong to the same identity.
 If the distance falls below a predefined threshold, the two images are recognized as belonging to the same identity; otherwise, they are classified as different identities.
-Based the measurement of FR, there exists the samples that can be classified as two different identities in theory.
-We denote these samples as multi-identity samples.
-Our experiments have verified the existence of multi-identity samples on common-used face dataset (refer to Appendix A).
-In impersonation attacks, the objective is to deceive the FR model into recognizing adversarial face examples as a predefined identity, which requires that the distance between the embeddings of the adversarial face example and the victim image be lower than the threshold.
-On the other hand, in dodging attacks, the aim is to trick the FR model into recognizing the adversarial face example as identities other than the attacker, which means that the distance between the embeddings of the adversarial face example and the attacker image should be greater than the threshold.
-The existence of multi-identity samples implies that a successful impersonation attack on FR does not necessarily guarantee a successful dodging attack on FR in theory.
-Our experiment confirms that the majority of adversarial face examples crafted by traditional impersonation attacks, which can successfully perform impersonation attacks, fail to perform dodging attacks in the black-box setting (see Appendix B).
-In practical application scenarios, criminals tend to generate adversarial examples with their own faces to deceive FR models into recognizing them as victims for impersonation attacks, while avoiding being recognized as criminals themselves for dodging attacks to escape legal prosecution.
-This necessitates the generation of adversarial examples that can perform both impersonation and dodging attacks.
-To improve the dodging performance of the adversarial face examples without degrading the impersonation performance, we propose a novel attack method, termed as Pre-training Pruning Restoration Attack (PPR).
-Specifically, inspired by the pruning methods in model compression, we prune the pre-trained adversarial face examples to sparsify the adversarial face examples.
-Subsequently, we add adversarial perturbations that favor dodging attacks on the pruned regions to improve the dodging performance of the adversarial face examples (i.e.
-the operation of Restoration).
-After pruning and restoration, the impersonation performance of the adversarial face examples can be maintained and the dodging performance of the adversarial face examples can be improved.
-Figure 1 illustrates a comparison between our proposed attack and the traditional impersonation attack.
-The adversarial face example generated by our proposed attack method successfully evades identification as the attacker in contrast with the adversarial face example crafted by the traditional impersonation attack method is identified as the attacker.
-The adversarial face example generated by our proposed attack method successfully evades identification as the attacker, while the adversarial face example crafted by the traditional impersonation attack method is identified as the attacker.
-Note that both the adversarial face examples crafted by our proposed attack method and the traditional impersonation attack method are capable of being identified as the victim image.
+Based on the measurement of FR, there are two decision boundaries for each FR model when crafting adversarial examples as shown in Fig.
+1.
+As a result, there exists the natural samples that can be classified as two different identities in theory.
+We denote these samples as multi-identity samples (refer to Section 4.4).
+The existence of multi-identity samples implies that a successful impersonation attack on FR does not necessarily guarantee a successful dodging attack on FR.
+Existing research indicates that an adversarial sample was located near the decision boundary [18; 4].
+Suppose we generate adversarial face examples using previous methods.
+In the white-box setting, both the structures and parameters of the victim models are known, enabling the generation of adversarial face examples that can cross the decision boundaries of both attacker and victim, as shown in Fig.
+1.
+However, in the black-box setting, the decision boundaries of black-box models differ from those of the surrogate models.
+Consequently, adversarial examples generated on the surrogate model lie near the decision boundary of the victim, preventing them from crossing the decision boundary of the attacker.
+As such, the majority of adversarial face examples crafted by previous methods, which can successfully perform impersonation attacks, fail to perform dodging attacks in the black-box setting.
+Figure 2: Overview of our Adv-Pruning attack framework, which consists of Priming, Pruning, and Restoration stages.
+(a) During the Priming stage, we optimize the adversarial examples to ensure they have sufficient attack performance.
+(b) In the Pruning stage, we propose Adversarial Priority Quantification to quantify the priority of adversarial perturbations.
+Subsequently, we sparsify the adversarial perturbations based on the quantified priorities.
+(c) In the Restoration stage, we present Biased Gradient Adaptation to introduces gradient perturbations biased to dodging attacks on the sparsified regions.
+In this paper, we propose a novel attack method, termed as Adversarial Pruning (Adv-Pruning).
+In the realm of adversarial attacks on FR, previous impersonation methods have achieved a significant level of sophistication.
+However, there remains a pressing need to bolster the efficacy of adversarial face examples in dodging attacks.
+Consequently, our research is directed towards enhancing the dodging attack performance of adversarial face examples while maintaining the impersonation attack performance.
+Specifically, we introduce an attack consisting of three stages: Priming, Pruning, and Restoration.
+In the Priming stage, we optimize the adversarial examples to ensure adequate attack potential.
+In the Pruning stage, with considering the pruning concept in model compression, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations using an priority measure which is directly proportional to the supremum of the absolute model output variances.
+After processing by Adversarial Priority Quantification, we prune the adversarial face examples to free up less prioritized adversarial perturbations.
+In the Restoration stage, we propose Biased Gradient Adaptation to add biased gradient perturbations favoring dodging attacks on the pruned regions to adapt the adversarial face examples into the space that can be classified as the victim while remaining unidentifiable as the attacker, thereby enhancing the dodging performance of the adversarial face examples without compromising the prioritized features of original adversarial perturbations.
+As illustrated in the top of Fig.
+1, after undergoing these stages, the adversarial face example generated by our proposed method can successfully traverse the decision boundaries of both the attacker and victim of the black-box model, achieving successful black-box impersonation and dodging attacks.
 Our main contributions are summarized as follows:
-  • We verify the universality of multi-identity samples among adversarial face examples crafted by traditional impersonation attacks and demonstrate that a successful impersonation attack does not necessarily equate to a successful dodging attack on FR systems.
-  • We propose a novel adversarial attack method called PPR based on adversarial example pruning.
-By pruning the adversarial face examples and subsequently introducing adversarial perturbations favoring dodging attacks, we effectively enhance the dodging performance.
-The approach allows us to maintain the impersonation performance while significantly improving the dodging performance of the adversarial face examples.
-  • Extensive experiments demonstrate that our proposed method achieves superior performance compared to the baseline adversarial attack methods.
-2 Related Works
+  • We offer a new perspective for adversarial attacks on FR models that the generation of adversarial examples should perform both impersonation and dodging attacks simultaneously.
+To the best of our knowledge, this is the first work that studies the universality of multi-identity samples among adversarial face examples crafted by impersonation attacks.
+  • We propose a novel adversarial attack method called Adversarial Pruning (Adv-Pruning).
+Adversarial Priority Quantification is presented to quantify the priority of the adversarial perturbations with minimal impact on absolute model output variances.
+Biased Gradient Adaptation is designed to adapt the adversarial examples to traverse both the decision boundaries of attacker and victim using biased gradients.
+  • Extensive experiments demonstrate that our proposed method achieves superior performance compared to the state-of-the-art adversarial attack methods.
+Moreover, our presented method could be plugged into various FR systems and adversarial attack methods.
+2 Related Work
 2.1 Adversarial Attacks
-The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images for the purpose of deceiving machine learning systems and causing them to make mistakes(Szegedy et al., 2014)(Goodfellow et al., 2015).
+The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images to deceive machine learning systems and cause them to make mistakes [47; 14].
 The existence of adversarial examples poses a significant threat to the security of current machine learning systems.
-Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems(Long et al., 2022)(Zhang et al., 2022)(Liang et al., 2023)(Wei et al., 2023)(Lu et al., 2023).
-To improve the performance of black-box adversarial attacks, DI is proposed by (Xie et al., 2019) to randomly transform adversarial examples in each iteration.
-VMI-FGSM, proposed by (Wang & He, 2021), employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance.
-SSA, as proposed by (Long et al., 2022), transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them.
-(Wang et al., 2023b) propose SIA.
-SIA applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculations.
-(Wang et al., 2023a) propose BSR.
-BSR divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation.
-2.2 Face Recognition
-FR is an open-set task, meaning we cannot predict the class of a user upon deploying an FR model.
-As a result, relying solely on the maximum probability value of the output of the neural network to determine the class the of a user becomes impractical.
-Instead, when deploying the FR model, we extract the embeddings of two images and compute the similarity between them.
-If the similarity exceeds a predefined threshold, we conclude that the two images belong to the same identity; otherwise, they belong to different identities.
-Recent research on FR primarily focuses on reducing the intra-class distance and increasing the inter-class distance of embeddings extracted by FR models.
-(Schroff et al., 2015) introduce FaceNet that employs triplet loss to map face images into a space where the distance represents the dissimilarity between them.
-(Wang et al., 2018) develop CosFace, which utilizes the large-margin cosine loss to enhance FR model performance.
-Additionally, (Deng et al., 2022) introduce ArcFace that incorporates the additive angular margin loss to train FR models that extract more discriminative features.
+Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems [33; 69; 30; 55; 34; 43; 12; 73; 68; 37].
+To improve the performance of black-box adversarial attacks, DI [60] applies random transformations to adversarial examples in each iteration to achieve a data augmentation effect.
+VMI-FGSM [53] employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance.
+SSA [33] transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them.
+SIA [54] applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculation.
+BSR [51] divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation.
+DA [16] utilizes dispersion amplification to enhance the multi-task attack capability of adversarial attacks.
+Despite their gratifying progress, these studies neglect the consideration of pruning adversarial examples through introducing pruning methods into the realm of adversarial attacks.
+In our research, we propose a novel pruning method capable of identifying and freeing up the adversarial perturbations with minimal impact on absolute model output variances, thereby sparsifying regions for adding adversarial perturbations with the aim of dodging capabilities improvement.
 2.2 Adversarial Attacks on Face Recognition
-Adversarial attacks on FR can be mainly classified into two categories: restricted attacks and unrestricted attacks.
+Based on the restriction of the adversarial perturbations, adversarial attacks on FR can be classified into two categories: restricted attacks [32; 61; 10; 74; 35; 31; 5] and unrestricted attacks [3; 6; 65; 8; 44; 46; 48; 56; 57].
 Restricted attacks on FR are the attacks that generate adversarial examples in a restricted bound (e.g.
 $L_{p}$ bound).
-Because FR belongs to the open-set task, the embedding-level loss is more suitable.
-Therefore, (Zhong & Deng, 2019) proposed FIM that use embedding-level loss to craft the adversarial face examples.
-(Zhong & Deng, 2021) proposed DFANet that performs dropout on the feature maps of the convolutional layers to improve the transferability.
-(Yang et al., 2021) propose TIP-IM to improve the face encryption.
-TIP-IM use MMD(Borgwardt et al., 2006) loss to improve the quality of encrypted face images and use the greedy insertion to select the optimal victim images.
-(Zhou et al., 2023b) propose BPFA to improve the transferability of adversarial attacks on FR by adding beneficial perturbations(Wen & Itti, ) on the features maps of the FR models to obtain effect that is similar to hard samples augmentation.
-The unrestricted adversarial attacks on FR are the attacks that generate adversarial examples without the restriction of a predefined bound.
-The unrestricted adversarial attacks on FR mainly focus on physical attacks(Xiao et al., 2021)(Yang et al., 2023)(Li et al., 2023a), attribute editing(Jia et al., 2022) and generating adversarial examples based on makeup transfer(Yin et al., 2021)(Hu et al., 2022)(Shamshad et al., 2023).
+To enhance the transferability of adversarial attacks on FR, [70] propose DFANet, which applies dropout on the feature maps of the convolutional layers to achieve ensemble-like effects.
+In addition, [72] introduces BPFA, which further improves the transferability of adversarial attacks on FR by incorporating beneficial perturbations [58] on the feature maps of the FR models, resulting in hard model augmentation effects.
+[29] leverages extra information from FR-correlated tasks and uses a multi-task optimization framework to enhance the transferability of crafted adversarial examples.
+The unrestricted adversarial attacks on FR are the attacks that generate adversarial examples without the restriction of a predefined perturbation bound.
+They mainly focus on physical attacks [59; 63; 28], attribute editing [40; 23] and generating adversarial examples based on makeup transfer [64; 19; 42].
+The existing literature on both restricted and unrestricted adversarial attacks on FR systems has successfully enhanced the performance of these attacks.
+Nevertheless, it remains under-explored in the correlation between impersonation and dodging attacks.
+This paper elegantly addresses this by investigating the correlation between impersonation and dodging attacks and introducing a novel attack method that bolsters the dodging capabilities while preserving the impersonation capabilities of previous methods.
 3 Methodology
 3.1 Problem Formulation
-In the context of our study, $\mathcal{F}^{vct}(x)$ denotes the FR model used by the victim to extract the embedding from a face image $x$.
+Let $\mathcal{F}^{vct}(x)$ denote the FR model used by the victim to extract the embedding from a face image $x$.
 We refer to $x^{s}$ and $x^{t}$ as the attacker and victim images, respectively.
-The objective of the impersonation attacks explored in our research is to manipulate $\mathcal{F}^{vct}$ into misclassifying $x^{adv}$ as $x^{t}$, while ensuring that $x^{adv}$ bears a close visual resemblance to $x^{s}$.
-To be more specific, the objective can be expressed as follows:
-$\begin{gathered}x^{adv}=\mathop{\arg\min}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (1)
-where $\mathcal{D}$ refers to a predefined distance metric, while $\epsilon$ specifies the maximum magnitude of permissible perturbation.
+The objective of the impersonation attacks explored in our research is to manipulate $\mathcal{F}^{vct}$ in order to misclassify $x^{adv}$ as $x^{t}$, while ensuring that $x^{adv}$ bears a close visual resemblance to $x^{s}$.
 By contrast, the objective of the dodging attacks proposed in this study is to render $\mathcal{F}^{vct}(x)$ unable to identify $x^{adv}$ as $x^{s}$, while simultaneously ensuring that $x^{adv}$ bears a visual resemblance to $x^{s}$.
-In a manner similar to impersonation attacks, the objective of dodging attacks can be formulated as follows:
-$\begin{gathered}x^{adv}=\mathop{\arg\max}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (2)
-Due to the differing objectives of the impersonation and dodging attacks, the calculation methods for the ASR also vary between the two attack types.
-When it comes to impersonation attacks on FR, the ASR can be computed as:
-${\rm ASR}^{i}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)<t^{i}\right)}{N_{p}}$ (3)
-where $N_{p}$ refers to the total number of face pairs and $t^{i}$ represents the impersonation attack threshold. For dodging attacks on FR systems, the ASR can be computed via the following formula:
-${\rm ASR}^{d}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)>t^{d}\right)}{N_{p}}$ (4)
-where the value of $t^{d}$ represents the threshold for dodging attacks.
+For the sake of clarity and conciseness, the detailed optimization objectives for both impersonation and dodging attacks are provided in the appendices.
+Few works explore the correlation between impersonation and dodging attacks on FR.
+In the following, we delve into the correlation between these two types of attacks and propose a novel method to enhance dodging attacks while maintaining impersonation attacks.
+An overview of the proposed method is illustrated in Fig.
+2.
+As depicted in Fig.
+2, our proposed method is structured into three stages: Priming, Pruning, and Restoration.
+Through the sequential application of these stages, we are able to generate adversarial examples that exhibit a potent combination of impersonation and dodging attack capabilities.
 3.2 Exploring the Impersonation and Dodging Attack on Face Recognition
-In most cases, the victim model $\mathcal{F}^{vct}$ is not accessible to the attacker, making it impossible to optimize objectives as stated in Equation 1 and Equation 2 directly.
-To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversaries that can be transferred to the victim model for an effective attack(Yuan et al., 2022)(Zhang et al., 2022)(Naseer et al., 2023)(Wang et al., 2023b)(Wang et al., 2023a).
+In most cases, the victim model $\mathcal{F}^{vct}$ is not accessible to the attacker, making it extremely challenging to optimize the objectives for black-box attacks directly.
+To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversarial examples that can be transferred to the victim model for an effective attack [66; 69; 39; 54; 51; 38; 11; 13; 27; 5].
 For impersonation attacks, the loss can be formulated as follows:
@@
 $x^{adv}$ is the adversarial example which is initialized with the same value as $x^{s}$.
-Dodging attacks can be formulated with the following loss function:
+The loss function of dodging attacks can be formulated as:
 $\mathcal{L}^{d}=-\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{s}\right)\right)\|^{2}_{2}$ (2)
-As the FR task is an open-set task, it is not possible to predict the classes of users during the practical deployment of the FR model.
-Therefore, we need to compare the similarity between two face images in order to discern whether they depict the same identity or not.
-Based on the identification method in face recognition, multi-identity samples exist theoretically.
+As the FR task is an open-set task, it is impractical to predict the classes of users during the practical deployment of the FR model.
+Therefore, we need to compare the distance between two face images to discern whether they depict the same identity or not.
+Based on the identification method in FR, multi-identity samples exist theoretically.
 Our experiments verify the existence of such samples among benign face images.
 The existence of multi-identity samples raises a question:
-To this end, we generated face adversarial samples using the traditional impersonation attack and evaluate its dodging ASR (refer to Appendix B).
-Our experiments demonstrate that the majority of adversarial face examples crafted by traditional impersonation attack cannot achieve a successful dodging attack in the black-box setting.
-Nonetheless, in real-world adversarial attacks, attackers do not want the adversarial face examples to be recognized as themselves, as this could lead to legal consequences.
+Does the success of an impersonation attack imply the success of dodging attacks on FR systems?
+To this end, we generate adversarial face examples using the previous impersonation attack and evaluate its dodging Attack Success Rate (ASR).
+Our experiment confirms that the majority of adversarial examples crafted through previous methods, which are successful in performing impersonation attacks, fail to successfully execute dodging attacks in the black-box setting (see Section 4.4).
+Nonetheless, in real-world adversarial attacks, attackers do not want the adversarial face examples to be recognized as themselves, as this may lead to legal consequences.
 Hence, it is crucial to research attack techniques that can execute both impersonation and dodging attacks simultaneously.
-Traditional impersonation attacks on FR systems have shown a remarkably high level of impersonation ASR in black-box settings.
-Therefore, our objective is to enhance the dodging performance while maintaining the impersonation effectiveness of traditional impersonation attacks.
-To accomplish this objective, a straightforward approach is to generate adversarial face examples by employing a Lagrangian attack strategy that utilizes the following loss function:
+Previous methods on FR systems have shown a remarkably high level of impersonation ASR in black-box settings.
+Therefore, our objective is to enhance the dodging performance while maintaining the impersonation effectiveness of previous attack methods.
+To accomplish this objective, a straightforward approach is to generate adversarial face examples using a multi-task attack strategy.
+In the following, we will take the Lagrangian attack strategy as the example for its simplicity.
+The Lagrangian attack strategy utilizes the following loss function to craft adversarial examples:
 $\mathcal{L}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$ (3)
-However, due to the conflict between the optimization between $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, there exists a trade-off between the performance of impersonation and dodging performance leading to subpar performance (See Section 4.2).
-If we alliviete the trade-off, we can achieve a better dodging performance while maintaining the impersonation performance.
-3.3 Pre-training Pruning Restoration Attack
+However, due to the conflict between the optimization between $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, there exists a trade-off between the performance of impersonation and dodging performance, leading to subpar performance (See Section 4.4).
+Suppose we can mitigate the trade-off, we will achieve a better dodging performance while maintaining the impersonation performance.
+3.3 Adversarial Pruning Attack
 To accomplish this objective, a straightforward approach is to fine-tune the adversarial face examples generated by the Lagrangian attack with a lower $\lambda$ value in order to enhance the performance of dodging attacks.
 However, this method does not enhance the dodging attack performance without compromising the impersonation attack performance (see Fine-tuning in Table 3).
-We contend that this issue arises because the newly introduced adversarial perturbation that favors dodging attacks ends up disrupting the pre-existing adversarial perturbation patterns.
+We contend that this issue arises because the newly introduced adversarial perturbation that favors dodging attacks ends up disrupting the prioritized features of existing adversarial perturbation.
 While it may improve the performance of dodging attacks, it inevitably diminishes the performance of impersonation attacks.
-To address this, we introduce new adversarial perturbations favoring dodging attacks in regions where the original perturbations have not been added.
-However, identifying suitable areas for these new perturbations is challenging due to their scarcity.
-Therefore, drawing inspiration from prior works in model pruning(Ma et al., 2023)(Frantar & Alistarh, 2023), we propose incorporating pruning into the realm of adversarial face examples.
-The aim is to free up some model parameters while minimizing the decrease in the performance of impersonation attacks, thereby creating space to introduce perturbations that facilitate dodging attacks.
-In order to prune the adversarial perturbation, our initial step is to assess its importance by identifying and sparsifying the less important regions.
-To estimate this importance, we propose utilizing the magnitude of the adversarial perturbation as a measure.
-A lower magnitude implies a lesser impact on the performance of the adversarial examples generated after sparsification.
-Let the adversarial examples be $x^{adv}$. The formula to calculate the importance can be express as:
+To address this, we introduce new perturbations favoring dodging attacks in regions where original perturbations are not added.
+Nevertheless, identifying suitable areas for these new perturbations is challenging due to their scarcity.
+Therefore, we propose a novel pruning method to release less prioritized adversarial perturbations with minimal impact on the absolute model output variances, thereby creating space to introduce perturbations that facilitate dodging attacks.
+Our proposed Adv-Pruning and be combined with various adversarial attacks.
+In the following, we will introduce our proposed Adv-Pruning based on Lagrangian attack in detail.
+In the Priming Stage, we utilize Eq.
+3 as the Priming loss $\mathcal{L}^{p}$ to craft the adversarial face examples:
+$x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)\right)$ (4)
+where $t$ is the iteration of the optimization process of adversarial examples, and $\beta$ is the step size when optimizing the adversarial face examples in the Priming stage, and $\prod\left(x\right)$ is the projection function that projects $x$ onto the $L_{p}$ norm bound.
+Adversarial Priority Quantification.
+After completing the Priming Stage, we obtain an adversarial example with varying magnitudes of gradient perturbations across different regions.
+Following this, we proceed to the Pruning stage to process the crafted adversarial example.
+In order to prune the adversarial perturbation, our initial step is to assess its priority.
+To estimate this priority, we propose Adversarial Priority Quantification to quantify the priority of adversarial perturbations.
+Specifically, Adversarial Priority Quantification utilizes the magnitude of the adversarial perturbation as a measure.
+A lower magnitude implies a lesser impact on the performance of the adversarial examples generated after sparsification, as the supremum of absolute model output variances is directly proportional to the magnitude of the adversarial perturbations.
+The proof is in the appendices.
+Let the adversarial examples be $x^{adv}$. The formula to calculate the priority can be expressed as:
 $\mathcal{I}=|x^{adv}-x^{s}|$ (5)
-where $\mathcal{I}\in\mathbb{R}^{\rm CHW}$. ${\rm C}$, ${\rm H}$, and ${\rm W}$ are the channel number, hight, width of the face images, respectively.
-Let $\kappa$ be the sparsity ratio for pruning the adversarial face examples that measures the ratio of adversarial perturbations to be set into zero.
-Let $s$ be the number of adversarial perturbation elements:
-$s={\rm CHW}$ (9)
-We arrange the elements in flattened vector of $\mathcal{I}$ in ascending order (from the lowest to the highest):
+where $\mathcal{I}\in\mathbb{R}^{\rm CHW}$. ${\rm C}$, ${\rm H}$, and ${\rm W}$ are the channel number, height, and width of the face images, respectively.
+Once the priority values of the adversarial perturbations are quantified, we employ these values to release less prioritized adversarial perturbations.
+Let $\kappa$ be the sparsity ratio for pruning the adversarial face examples that measure the ratio of perturbations to be set into zero.
+Let $s={\rm CHW}$ be the number of adversarial perturbation elements.
+We arrange the elements in a flattened vector of $\mathcal{I}$ in ascending order (from the lowest to the highest):
 $\mathcal{Q}={\rm Sort}\left(\Psi\left(\mathcal{I}\right)\right)$ (6)
 where $\Psi$ is the flatten operation.
-Let $\mathcal{W}$ be the set of the elements of the adversarial perturbations to be pruned.
-Given the importance calculation method for pruning, the value of $\mathcal{W}$ can be calculated as follows:
+Let $\mathcal{W}$ be the set of the elements of the adversarial perturbations to be pruned. Given the priority calculation method for pruning, the value of $\mathcal{W}$ can be calculated as follows:
 $\mathcal{W}=\mathcal{Q}\left[:\kappa s\right]$ (7)
-where the colon denotes the slice operation to obtain the first $\kappa s$ elements.
-The mask for pruning can be obtained using $\mathcal{W}$:
+where the colon denotes the slice operation to obtain the first $\kappa s$ elements. The pruning mask, which has the same shape as $\mathcal{I}$, can be obtained by utilizing $\mathcal{W}$:
 $\mathcal{M}_{i,j,k}=\begin{cases}0,&\text{if }\mathcal{I}_{i,j,k}\in\mathcal{W}\\ 1,&\text{if }\mathcal{I}_{i,j,k}\notin\mathcal{W}\end{cases}$ (8)
@@
 $\bar{x}^{adv}=x^{s}+\left(x^{adv}-x^{s}\right)\odot\mathcal{M}$ (9)
-Once the adversarial face examples are pruned, we can proceed to restore the adversarial perturbations in the pruned regions.
+where $\bar{x}^{adv}$ is the adversarial face example after pruning.
+Biased Gradient Adaptation.
+During the Restoration stage, we restore the adversarial face examples in the previously pruned region using our proposed Biased Gradient Adaptation.
+Biased Gradient Adaptation using the following loss function to craft gradient biased to the dodging attacks to adapt the crafted adversarial examples into the space that favors dodging attacks.
+$\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$ (10)
+where $\tilde{\lambda}$ is a weight that is lower than $\lambda$ that is objective for crafting adversarial face examples that favor dodging attacks.
 The mask representing the regions for restoring the adversarial examples can be denoted as:
 $\mathcal{A}=1-\mathcal{M}$ (11)
-For both the pre-training and restoration stages, various adversarial attacks can be employed.
-In the following, we will illustrate the pre-training and restoration stages of of our proposed method by using the Lagrangian attack as an example.
-In the stage of pre-training, we utilize Equation 7 as the pre-training loss $\mathcal{L}^{p}$ to craft the adversarial face examples:
-$x^{adv}_{t}=x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)$ (15)
-where $t$ is the iteration of optimization process of adversarial examples, $\beta$ is the step size when optimizing the adversarial face examples in the pre-training stage.
-Once the adversarial examples are pre-trained, we utilize our proposed method of pruning adversarial face examples to free up some adversarial perturbations.
-During the restoration stage, we restore the adversarial face examples in the previously pruned region using the following loss function:
-$\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$ (16)
-where $\tilde{\lambda}$ is a weight which is lower than $\lambda$ that is objective for crafting adversarial face examples that favor dodging attacks.
-$x^{adv}_{t}=x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)$ (17)
-where $\gamma$ is the step size when optimizing the adversarial face examples in the restoration stage. The algorithm of our proposed method based on Lagrangian attack are illustrated in Algorithm 1.
-Algorithm 1 Pre-training Pruning Restoration Attack Based on Lagrangian
-0:  Negative face image pair $\{x^{s},x^{t}\}$, the step size of the adversarial perturbations $\beta$, the number of the iterations of pre-training stage $n$, the maximum number of iterations $m$, maximum allowable perturbation magnitude $\epsilon$, the surrogate FR model $\mathcal{F}$.
+Subsequently, we utilize the following formula to restore the pruned adversarial face examples:
+$\begin{split}x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{n}+\mathcal{A}\odot\left(x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)-\bar{x}^{adv}\right)\right)\end{split}$ (12)
+where $\gamma$ is the step size when optimizing the adversarial face examples in the Restoration stage, $\bar{x}^{adv}$ is the adversarial example crafted by the Priming Stage, and $\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}$ is the biased gradient.
+The pseudo-code of our proposed method based on the Lagrangian attack is illustrated in the appendices.
+4 Experiments
+4.1 Experimental Setting
+Datasets.
+We opt to use the LFW [20], CelebA-HQ [24], and FFHQ [25] datasets for our experiments.
+LFW serves as an unconstrained face dataset for FR.
+CelebA-HQ and FFHQ consists of high-quality images.
+The LFW and CelebA-HQ utilized in our experiments are identical to those employed in [72; 71], while FFHQ is the corresponding dataset provided by the Sibling-Attack official page, ensuring the consistency for analysis.
+Figure 3: The ASR on FR models trained by multiple algorithms.
+Figure 4: Comparisons of ASR (%) on LFW with adversarial robust models as victim models.
+Table 1: Comparisons of dodging ASR (%) results for attacks on the LFW and CelebA-HQ datasets.
+The surrogate models are presented in the first column, and the victim models are listed in the second row.
+The numbers before and after the slash represent the results of the baseline attack and the attack that combines the baseline attack with our proposed attack, respectively.
+LFW CelebA-HQ
+Surrogate Model Attack IR152 FaceNet MF IR152 FaceNet MF
+IR152 [17] DI 95.4 / 100.0 5.8 / 11.3 0.2 / 0.8 87.9 / 100.0 8.4 / 16.4 0.4 / 1.2
+VMI 92.7 / 100.0 17.3 / 32.5 1.2 / 9.5 92.2 / 100.0 14.3 / 27.9 1.2 / 4.1
+SSA 78.8 / 100.0 5.7 / 22.5 0.9 / 10.6 83.8 / 99.9 7.2 / 19.6 0.4 / 5.6
+DFANet 98.9 / 100.0 1.4 / 4.2 0.0 / 0.3 98.9 / 100.0 2.3 / 6.0 0.0 / 0.4
+SIA 81.7 / 100.0 13.0 / 37.5 0.8 / 8.9 78.4 / 100.0 13.2 / 35.6 0.7 / 7.4
+BSR 52.4 / 100.0 5.3 / 17.6 0.1 / 1.5 48.5 / 99.9 5.4 / 18.0 0.3 / 1.9
+BPFA 92.6 / 100.0 1.7 / 7.3 0.0 / 1.2 90.4 / 100.0 2.1 / 8.1 0.1 / 0.8
+FaceNet [41] DI 5.3 / 10.3 99.8 / 99.9 3.1 / 10.3 1.5 / 3.1 99.4 / 99.9 1.8 / 4.7
+VMI 9.7 / 14.3 99.8 / 99.9 6.2 / 13.2 3.1 / 7.1 99.3 / 99.8 3.6 / 9.3
+SSA 6.0 / 14.0 97.5 / 99.9 6.6 / 26.2 2.0 / 5.5 96.9 / 99.7 4.2 / 14.6
+DFANet 1.6 / 3.3 99.8 / 99.9 0.4 / 2.7 0.5 / 2.6 99.1 / 100.0 0.8 / 4.1
+SIA 11.2 / 20.6 99.5 / 99.9 8.7 / 21.2 4.0 / 8.9 99.4 / 99.9 5.4 / 13.7
+BSR 12.2 / 19.2 98.6 / 99.9 9.0 / 17.8 4.6 / 10.1 98.8 / 99.9 5.3 / 14.1
+BPFA 4.7 / 16.8 98.6 / 100.0 1.6 / 15.0 1.1 / 4.2 99.0 / 100.0 0.6 / 5.1
+MF [9] DI 2.2 / 7.3 18.2 / 36.4 99.2 / 100.0 0.1 / 2.5 12.1 / 31.3 95.2 / 100.0
+VMI 1.0 / 2.8 8.4 / 20.9 99.7 / 100.0 0.2 / 0.4 5.2 / 15.0 98.2 / 100.0
+SSA 0.7 / 4.1 6.1 / 23.5 98.3 / 100.0 0.0 / 0.6 3.9 / 18.5 93.3 / 100.0
+DFANet 0.2 / 1.0 1.5 / 5.8 99.6 / 100.0 0.0 / 0.2 1.1 / 7.7 99.1 / 100.0
+SIA 1.0 / 5.9 10.6 / 36.6 98.4 / 100.0 0.1 / 2.4 9.0 / 24.4 96.3 / 100.0
+BSR 0.4 / 1.5 3.7 / 14.7 84.9 / 100.0 0.1 / 0.6 2.9 / 12.6 77.6 / 100.0
+BPFA 0.9 / 4.1 4.6 / 20.4 97.7 / 100.0 0.0 / 2.3 4.0 / 20.4 96.2 / 100.0
+Face Recognition Models.
+The normal trained FR models employed in our experiments include IR152 [17], FaceNet [41], MobileFace (abbreviated as MF) [9], ArcFace [9], CircleLoss [45], CurricularFace [22], MagFace [36], MV-Softmax [52], and NPCFace [67].
+IR152, FaceNet, and MF are identical to those used in [64; 19; 72; 71].
+ArcFace, CircleLoss, CurricularFace, MagFace, MV-Softmax, and NPCFace are the official models available in FaceX-ZOO [50].
+Additionally, we incorporate adversarial robust FR models in our experiments, denoted as IR152adv, FaceNetadv, and MFadv, which are identical to those used in [72].
+For calculating the ASR in impersonation and dodging attacks, we choose the thresholds based on FAR@0.001 on the entire LFW dataset.
+Attack Setting.
+Without any particular emphasis, we set the maximum allowable perturbation magnitude to 10 based on the $L_{\infty}$ norm bound and utilize the Lagrangian attack method as the attack in both the Priming and Restoration stages.
+Additionally, we specify the maximum number of iterative steps as 200.
+For both the Priming and Restoration stages, the step size is uniformly designated as 1.0.
+Evaluation Metrics.
+We employ Attack Success Rate (ASR) to evaluate the performance of various attacks.
+ASR signifies the proportion of successfully attacked adversarial examples out of all the adversarial examples.
+We use ASRi and ASRd to denote impersonation and dodging ASR, respectively.
+The detailed calculation methods for ASRi and ASRd are provided in the appendices.
+Compared methods.
+Our proposed attack is a restricted attack method that aims to maliciously attack FR systems to expose more blind spots of them.
+It is not fair to compare our proposed method with unrestricted attacks that do not limit the magnitude of the adversarial perturbations.
+Therefore, we choose restricted attacks on FR that aim to maliciously attack FR systems [70] [72] [29] and state-of-the-art transfer attacks [60] [33] [54] [51] as our baseline.
+Table 2: Comparisons of ASR (%) with multi-task attacks on LFW dataset. Models in the second row are victim models.
+ASRd ASRi
+Attack IR152 FaceNet MF
+Lagrangian 3.9 26.5 100.0 26.0
+Lagrangian + ours 7.3 36.4 100.0 26.6
+DA 11.0 35.6 99.1 37.4
+DA + ours 17.5 44.9 99.4 37.8
+4.2 Comparison Study
+We compare our proposed attack method with the state-of-the-art attacks on multiple FR models and datasets.
+Several adversarial examples are illustrated in Fig.
+5.
+The attack performance results are shown in Table 1.
+Table 1 illustrates that the incorporation of our proposed attack method significantly enhances the dodging ASR of adversarial attacks.
+It is worth noting that the average black-box impersonation ASRs of the baseline attacks in Table 1 also increase after integrating our proposed attack method.
+This demonstrates the effectiveness of our proposed method in improving the dodging attack performance while simultaneously maintaining the impersonation attack performance.
+Furthermore, we conducted a comparison between our proposed Adv-Pruning and multi-task attacks using MF as the surrogate model on LFW based on DI.
+For our proposed method, we choose the corresponding multi-task attack as the attack for both the Priming and Restoration stages.
+The dodging ASR and average black-box impersonation ASR results are shown in Table 2.
+Table 2 underscores the effectiveness of our method in enhancing the dodging performance of multi-task attacks while maintaining the impersonation performance.
+To further validate our proposed attack method on additional FR models, we selected SIA [54] as Baseline and IR152 as the surrogate model.
+The experimental settings are consistent with those described in Table 1.
+The dodging ASR across multiple FR models is demonstrated in Fig.
+3.
+As depicted in Fig.
+3, the dodging ASR improves on multiple FR models after integrating our proposed method, further confirming the effectiveness of our attack.
+Figure 5: The Illustration of adversarial examples crafted by various attacks.
+First column: some attacker images.
+Last column: the corresponding victim images.
+The second to fifth columns exhibit the corresponding adversarial face examples crafted by BPFA, BPFA + ours, BSR, and BSR + ours, respectively.
+In practical application scenarios, victims can employ adversarial robust models to defend against adversarial attacks.
+Consequently, it becomes crucial to evaluate the performance of adversarial attacks on these robust models.
+In this study, we generate adversarial examples on the LFW dataset using MF as the surrogate model and assess the performance of various attacks on the adversarial robust models.
+The results are presented in Fig.
+4.
+The letters following the en dash represent the surrogate models, with ’I’, ’F’, and ’M’ corresponding to IR152adv, FaceNetadv, and MFadv, respectively.
+Fig.
+4 illustrates that the inclusion of our proposed method leads to improvements in both dodging and impersonation performance.
+These results serve as evidence of the effectiveness of our proposed method on adversarial robust models.
+JPEG compression is a widely adopted method for image compression during transmission, concurrently acting as a defense mechanism against adversarial examples.
+To assess the effectiveness of our proposed attack under JPEG compression, we utilize DI as the baseline attack and MF as the surrogate model, evaluating the attack performance on ArcFace and CurricularFace models with experimental settings consistent with those described in Table 1.
+The results are illustrated in Fig.
+6.
+These results demonstrate that across varying levels of JPEG compression, our proposed attack method consistently outperforms the baseline attack, thereby highlighting its effectiveness under JPEG compression.
+The experimental results on negative cosine similarity loss, and Sibling-Attack are presented in the appendices.
+4.3 Ablation Study
+To delve into the properties of our proposed attack method, we conducted an ablation experiment using DI as the Baseline attack, with MF serving as the surrogate model on the LFW dataset.
+To confirm the effectiveness of our pruning method, we employed the Random Zeroing (RZ) method, which randomly sets adversarial perturbations to zero.
+We applied this method and our pruning method to free up 20% of the adversarial perturbations crafted by the Lagrangian attack.
+For Fine-tuning, we employed the Lagrangian attack as the method to further optimize the Lagrangian adversarial examples with a lower $\lambda$.
+The dodging attack ASR and average black-box impersonation ASR results are shown in Table 3.
+Table 3 demonstrates that our proposed pruning method for adversarial examples achieves a significantly smaller decrease in ASR than RZ after pruning 20% of adversarial perturbations, indicating the effectiveness of our pruning method.
+After being processed using the Pruning and Restoration stage of our proposed Adv-Pruning method, both impersonation and dodging ASR of the crafted adversarial face examples are recovered and higher than the Baseline attack method.
+These results demonstrate the effectiveness of our proposed Adv-Pruning in improving the dodging performance of adversarial attacks on FR without compromising the impersonation attack performance.
+The sparsity ratio quantifies the proportion of adversarial perturbations that are allowed to be discarded during the Pruning stage.
+This ratio greatly impacts the performance of our proposed attack method.
+Hence, we conducted a sensitivity study on the sparsity ratio to analyze its effect on performance of the algorithm.
+The Lagrangian attack method based on DI is selected as the Baseline.
+We conduct a hyperparameter sensitivity study on LFW using FaceNet as the surrogate model, and adjust the value of $\tilde{\lambda}$ to ensure that the average black-box impersonation ASR results were within a 0.4% absolute difference compared to the Baseline.
+The dodging ASR results are illustrated in the right plot of Fig.
+7.
+The results illustrate that the dodging ASR of our proposed method initially increases and then decreases as the sparsity ratio increases.
+When the sparsity ratio increases, a greater number of adversarial perturbations are pruned, creating more empty regions for the adversarial perturbations that favor dodging attacks in the Restoration stage.
+If the sparsity ratio is set to a too-high value, an excessive number of adversarial perturbations are allowed to be freed up, resulting in a degradation of performance for the adversarial examples crafted by the Priming stage.
+Consequently, the performance of adversarial face examples will decrease.
+Figure 6: The dodging ASR under various JPEG Q values.
+4.4 Analytical Study
+Multi-identity Samples among the Natural Face Images: Multi-identity samples are intriguing samples that can be classified as multiple classes in FR.
+In this section, we will explore the existence of multi-identity samples among the natural face images.
+Table 3: Comparisons of ASR (%) results of dodging attack and impersonation attacks on the LFW dataset. The models in the second row are the victim models.
+ASRd ASRi
+Attack IR152 FaceNet MF
+Baseline 2.2 18.2 99.2 25.6
+Lagrangian 3.9 26.5 100.0 26.0
+Fine-tuning 4.2 26.3 100.0 25.6
+RZ 0.6 4.4 94.8 15.1
+Pruning 3.4 24.8 100.0 25.4
+Adv-Pruning 5.4 32.5 100.0 26.3
+We randomly select negative face pairs from the entire LFW dataset.
+Subsequently, we use MF as our FR model to extract the embeddings of the face images in each face pair and calculate the cosine similarity between the two images.
+If the cosine similarity surpasses the threshold, both images in the pair are classified as belonging to multiple identities, indicating that they are multi-identity samples.
+Our findings demonstrate the presence of multi-identity samples among the benign face images, as illustrated in the bottom left of Fig.
+1.
+The multi-identity samples in Fig.
+1 closely resemble the appearances of the identities they are classified into.
+To analyze the cause of this phenomenon, we need to consider the properties of both the multi-identity samples and the FR model.
+Commonly-used FR models are well-trained and capable of correctly classifying the majority of benign face images.
+However, there are some benign face images that the FR model fails to classify accurately, and these samples are referred to as hard samples [21; 72].
+Multi-identity samples are a specific type of hard sample known as hard negative samples.
+Typically, hard negative samples exhibit a similar appearance [67], indicating that the multi-identity samples among the benign face images share a resemblance.
+Universality of Multi-identity Samples among Adversarial Face Examples: The previous impersonation methods craft adversarial face examples only use the impersonation loss $\mathcal{L}^{i}$.
+In this section, we will investigate the ratio of multi-identity samples and evaluate the effectiveness of previous impersonation methods in terms of dodging attacks.
+We utilize the Multi-identity Sample Ratio (MSR) to gauge the proportion of multi-identity samples in the adversarial face examples capable of executing successful impersonation attacks.
+These multi-identity samples can be recognized as both the attacker and victim identities in the setting of our paper.
+The detailed calculation method of MSR is in the appendices.
+We evaluate the MSR, impersonation ASR and dodging ASR using MF as the surrogate model on the LFW dataset and the results are demonstrated in the appendices.
+The results demonstrates that most of the crafted adversarial face examples are multi-identity samples in the black-box setting.
+This indicates that most adversarial face examples generated through previous impersonation attacks are unable to attain a successful dodging attack in the black-box setting.
+Nevertheless, in the white-box setting, the majority of adversarial face examples capable of executing successful impersonation attacks also demonstrate success in dodging attacks.
+To analyze the reason, we consider the metric used to determine whether two face images belong to the same identity in FR.
+Since FR is an open-set task, we rely on the distance in the embedding space to make decisions.
+The top of Fig.
+1 illustrates two decision boundaries for each FR model, one for attacker identity and one for victim identity.
+In the white-box setting, if we generate adversarial examples using $\mathcal{L}^{i}$, these adversarial examples can penetrate a space where they are recognized as the victim identity rather than the attacker identity.
+However, the decision boundary of the black-box model differs from that of the surrogate model.
+In the black-box setting, most adversarial examples are found between the decision boundaries of the black-box model, resulting in the majority of adversarial examples crafted using $\mathcal{L}^{i}$ being multi-identity samples.
+This demonstrates that the adversarial face examples are positioned near the decision boundary in the black-box setting.
+The Trade-off Between the Impersonation Attacks and Dodging Attacks: Owing to the inherent conflict during the optimization process of impersonation and dodging losses in the black-box setting, there exists a trade-off between impersonation and dodging performance.
+We craft adversarial face examples using Lagrangian attack and our proposed Adv-Pruning on LFW based on DI.
+The average black-box results are demonstrated in the left plot of Fig.
+7.
+Figure 7: (a) The trade-off between the ASR (%) of impersonation attack and dodging attack of adversarial examples. (b) The dodging ASR (%) in different sparsity ratios.
+The results illustrate that our proposed method can reduce the trade-off between impersonation and dodging performance in the black-box setting.
+The pruning operation of our proposed Adv-Pruning serves to sparsify the adversarial perturbations while preserving the impersonation performance.
+On the other hand, the restoration operation tends to introduce adversarial perturbations in the pruned areas, specifically favoring dodging attacks.
+These operations effectively enhance the dodging attack performance while maintaining the impersonation attack performance, ultimately mitigating the trade-off.
+5 Conclusion
+In this paper, we reveal the universality of multi-identity samples among adversarial face examples crafted by previous impersonation attacks and the success of an impersonation attack may not necessarily imply the success of dodging attacks on FR systems in the black-box setting.
+In order to improve dodging performance without compromising impersonation performance, we proposed a novel attack, namely Adv-Pruning.
+Adv-Pruning comprises Priming, Pruning, and Restoration Stages.
+Leveraging our proposed Adversarial Priority Quantification, we identify less prioritized adversarial perturbations with minimal impact on absolute model output variances.
+Through our proposed Biased Gradient Adaptation, biased gradient perturbations are applied to the sparsified regions, adapting adversarial face examples to a space favoring dodging attacks.
+Extensive experiments demonstrate the effectiveness of our proposed method.
+Appendix A The Pseudo-code of Adversarial Pruning Attack Method
+The pseudo-code of our proposed method based on the Lagrangian attack is illustrated in Algorithm 1.
+Algorithm 1 Adversarial Pruning Attack Based on Lagrangian
+0:  Negative face image pair $\{x^{s},x^{t}\}$, the number of the iterations of Priming stage $n$, the maximum number of iterations $m$, maximum allowable perturbation magnitude $\epsilon$, the surrogate FR model $\mathcal{F}$, the step size $\beta$ and $\gamma$.
 0:  An adversarial face example $x^{adv}_{m}$
@@
 2:  for $t=1,...,n$ do
-3:     $\mathcal{L}^{p}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Pre-training Stage
+3:     $\mathcal{L}^{p}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Priming Stage
 4:     $x^{adv}_{t}=x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)$
@@
 9:  $\mathcal{W}=\mathcal{Q}\left[:\kappa s\right]$
-10:  Get $\mathcal{M}$ by Equation 12.
+10:  Get $\mathcal{M}$ by Eq. 8.
 11:  $x^{adv}_{n}=x^{s}+\left(x^{adv}_{n}-x^{s}\right)\odot\mathcal{M}$
@@
 18:  end for
-4 Experiments
-4.1 Experimental Setting
-Datasets.
-We have opted to use the LFW dataset (Huang et al., 2007) and the CelebA-HQ dataset (Karras et al., 2018) for our experiments.
-LFW serves as an unconstrained face dataset for FR, while CelebA-HQ consists of high-quality images.
-These two datasets are widely employed in research on adversarial attacks in FR (Xiao et al., 2021; Yin et al., 2021; Hu et al., 2022).
-The LFW and CelebA-HQ datasets utilized in our experiments are identical to those employed in previous studies (Zhou et al., 2023b, a).
-Table 1: Comparisons of dodging ASR results for attacks on the LFW and CelebA-HQ datasets. The surrogate models are presented in the first column, and the victim models are listed in the second row.
-LFW CelebA-HQ
-Surrogate Model Attack IR152 FaceNet MF IR152 FaceNet MF
-IR152 DI 95.4 / 100.0 5.8 / 11.3 0.2 / 0.8 87.9 / 100.0 8.4 / 16.4 0.4 / 1.2
-VMI 92.7 / 100.0 17.3 / 32.5 1.2 / 9.5 92.2 / 100.0 14.3 / 27.9 1.2 / 4.1
-SSA 78.8 / 100.0 5.7 / 22.5 0.9 / 10.6 83.8 / 99.9 7.2 / 19.6 0.4 / 5.6
-DFANet 98.9 / 100.0 1.4 / 4.2 0.0 / 0.3 98.9 / 100.0 2.3 / 6.0 0.0 / 0.4
-SIA 81.7 / 100.0 13.0 / 37.5 0.8 / 8.9 78.4 / 100.0 13.2 / 35.6 0.7 / 7.4
-BSR 52.4 / 100.0 5.3 / 17.6 0.1 / 1.5 48.5 / 99.9 5.4 / 18.0 0.3 / 1.9
-BPFA 92.6 / 100.0 1.7 / 7.3 0.0 / 1.2 90.4 / 100.0 2.1 / 8.1 0.1 / 0.8
-FaceNet DI 5.3 / 10.3 99.8 / 99.9 3.1 / 10.3 1.5 / 3.1 99.4 / 99.9 1.8 / 4.7
-VMI 9.7 / 14.3 99.8 / 99.9 6.2 / 13.2 3.1 / 7.1 99.3 / 99.8 3.6 / 9.3
-SSA 6.0 / 14.0 97.5 / 99.9 6.6 / 26.2 2.0 / 5.5 96.9 / 99.7 4.2 / 14.6
-DFANet 1.6 / 3.3 99.8 / 99.9 0.4 / 2.7 0.5 / 2.6 99.1 / 100.0 0.8 / 4.1
-SIA 11.2 / 20.6 99.5 / 99.9 8.7 / 21.2 4.0 / 8.9 99.4 / 99.9 5.4 / 13.7
-BSR 12.2 / 19.2 98.6 / 99.9 9.0 / 17.8 4.6 / 10.1 98.8 / 99.9 5.3 / 14.1
-BPFA 4.7 / 16.8 98.6 / 100.0 1.6 / 15.0 1.1 / 4.2 99.0 / 100.0 0.6 / 5.1
-MF DI 2.2 / 7.3 18.2 / 36.4 99.2 / 100.0 0.1 / 2.5 12.1 / 31.3 95.2 / 100.0
-VMI 1.0 / 2.8 8.4 / 20.9 99.7 / 100.0 0.2 / 0.4 5.2 / 15.0 98.2 / 100.0
-SSA 0.7 / 4.1 6.1 / 23.5 98.3 / 100.0 0.0 / 0.6 3.9 / 18.5 93.3 / 100.0
-DFANet 0.2 / 1.0 1.5 / 5.8 99.6 / 100.0 0.0 / 0.2 1.1 / 7.7 99.1 / 100.0
-SIA 1.0 / 5.9 10.6 / 36.6 98.4 / 100.0 0.1 / 2.4 9.0 / 24.4 96.3 / 100.0
-BSR 0.4 / 1.5 3.7 / 14.7 84.9 / 100.0 0.1 / 0.6 2.9 / 12.6 77.6 / 100.0
-BPFA 0.9 / 4.1 4.6 / 20.4 97.7 / 100.0 0.0 / 2.3 4.0 / 20.4 96.2 / 100.0
+Appendix B The Optimization Objective of the Impersonation and Dodging Attacks
+The objective of impersonation attacks in our research is to craft an adversarial example $x^{adv}$ that causes the model $\mathcal{F}^{vct}$ to incorrectly classify it as the target sample $x^{t}$, while simultaneously maintaining a high degree of visual similarity between $x^{adv}$ and the original sample $x^{s}$.
+To be more specific, the objective can be expressed as follows:
+$\begin{gathered}x^{adv}=\mathop{\arg\min}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (13)
+where $\mathcal{D}$ refers to a predefined distance metric, while $\epsilon$ specifies the maximum magnitude of permissible perturbation.
+In contrast, the dodging attacks introduced in this study aim to prevent the model $\mathcal{F}^{vct}$ from correctly recognizing $x^{adv}$ as the source sample $x^{s}$.
+This is achieved while maintaining a visually similar appearance between the adversarial example $x^{adv}$ and the original input $x^{s}$.
+In a manner similar to impersonation attacks, the objective of dodging attacks can be formulated as follows:
+$\begin{gathered}x^{adv}=\mathop{\arg\max}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (14)
+Appendix C The proof of the supremum of the absolute variance
+In this section, we will provide the proof of the supremum of the absolute variance of the output of $\mathcal{F}$ before and after pruning. The absolute variance can be expressed as follows:
+$\rho=\|\mathcal{F}\left(x+\delta\right)-\mathcal{F}\left(x\right)\|$ (15)
 […]
-For calculating the attack success rate in impersonation and dodging attacks, we choose the thresholds based on FAR@0.001 on the entire LFW dataset.
-Attack Setting.
-Without any particular emphasis, we set the maximum allowable perturbation magnitude, denoted as $\epsilon$, to 10.
-This value is based on the $L_{\infty}$ norm bound, considering pixel values within the range of $[0,255]$.
-Furthermore, we specify the maximum number of iterative steps as 200.
-Moreover, both $\beta$ and $\gamma$ are set to 1.0.
-Evaluation Metrics. We employ the attack success rate (ASR) to assess the performance of various attacks. ASR represents the ratio of successfully attacked adversarial face examples.
-Baseline methods Our proposed attack is a restricted attack method that aims to enhance the robustness of FR models.
-It is not rational to compare our proposed method with the unrestricted attacks which do not limit the magnitude of the adversarial perturbations.
-Therefore, we choose restricted attacks on FR that aim to enhance the robustness of FR models(Zhong & Deng, 2021)(Zhou et al., 2023b) and transfer attacks(Xie et al., 2019)(Long et al., 2022)(Wang et al., 2023b)(Wang et al., 2023a) as our baseline.
-Table 2: Comparisons of dodging ASR with multi-task attacks on the LFW dataset. The models in the first row are the victim models.
-[…]
-IR152 FaceNet MF
-Lagrangian 3.9 26.5 100.0 26.0
-Lagrangian + ours 7.3 36.4 100.0 26.6
-DA 11.0 35.6 99.1 37.4
-DA + ours 17.5 44.9 99.4 37.8
-4.2 The Trade-off Between the Impersonation Attacks and Dodging Attacks
-Owing to the inherent conflict during the optimization process of impersonation and dodging losses, there exists a trade-off between impersonation and dodging performance.
-We craft adversarial face examples using Lagrangian attack and our proposed PPR on LFW based on DI.
-The average black-box results are demonstrated in Figure 2.
-Figure 2: The trade-off between the performance of impersonation attack and dodging attack of adversarial examples.
-Figure 2 illustrates that the inclusion of our proposed method into the Lagrangian attack results in reduced the trade-off.
-It is not hard to see the reason.
-The pruning operation of our proposed PPR serves to sparsify the adversarial perturbations while preserving the impersonation performance.
-On the other hand, the restoration operation tends to introduce adversarial perturbations in the pruned areas, specifically favoring dodging attacks.
-These operations effectively enhance the dodging attack performance while maintaining the impersonation attack performance, ultimately mitigating the trade-off.
-4.3 Comparison Study
-We compare our proposed attack method with the baseline attacks on multiple FR models and datasets.
-For our proposed method, we choose the Lagrangian attack as the attack for both the stages of pre-training and restoration.
-The results are shown in Table 1.
-The numbers before the slash represent the results of the baseline attack, which refers to the attack method mentioned in the seconds column that solely utilizes the traditional impersonation attacks to create the adversarial examples.
-On the other hand, the numbers after the slash correspond to the results obtained by combining the baseline attack with our proposed PPR.
-Table 1 shows that after being added our proposed attack method, the dodging ASR of the adversarial attacks improve a lot.
-Note that the average black-box impersonation ASRs of the baseline attacks in Table 1 increase after incorporating our proposed attack method.
-This demonstrate the effectiveness of our proposed method in improving the dodging attack performance while maintain the impersonation attack performance.
-[…]
-In practical application scenarios, victims can employ adversarial robust models to defend against adversarial attacks.
-Consequently, it becomes crucial to evaluate the performance of adversarial attacks on these robust models.
-In this study, we generate adversarial examples on the LFW dataset using MF as the surrogate model and assess the performance of various attacks on the adversarial robust models.
-The results are presented in Table 3.
-Table 3 illustrates that the inclusion of our proposed method leads to improvements in both dodging and impersonation performance.
-These results serve as evidence of the effectiveness of our proposed method on adversarial robust models.
-Table 3: Comparisons of dodging ASR on LFW with adversarial robust models as victim models.
-[…]
-[…]
-DI 2.3 7.8 0.5 19.8
-DI + ours 7.2 17.1 4.7 20.3
-VMI 0.8 4.5 0.3 15.8
-VMI + ours 3.3 8.9 3.0 16.2
-SSA 0.6 2.9 0.1 14.7
-SSA + ours 3.9 9.7 3.0 14.8
-DFANet 0.2 0.8 0 8.1
-DFANet + ours 0.3 1.3 0.1 8.3
-SIA 1.4 6 0.2 17.5
-SIA + ours 5.6 15.5 5.6 17.6
-BSR 0.4 1.9 0 6.4
-BSR + ours 0.7 3.5 0.4 6.6
-BPFA 0.3 1.8 0.1 13.1
-BPFA + ours 4.1 8.4 4.2 13.4
-4.4 Ablation Study
-To delve deeper into the properties of our proposed attack method, we conducted an ablation experiment using DI as the Baseline attack, with MF serving as the surrogate model on the LFW dataset.
-To confirm the effectiveness of our pruning method, we employed the Random Zeroing (RZ) method, which randomly sets adversarial perturbations to zero.
-We applied this method and our pruning method to free up 20% of the adversarial perturbations.
-For Fine-tuning, we employed the Lagrangian attack as the method to further optimize the adversarial examples crafted by the Lagrangian attack.
-The dodging attack ASR and average black-box impersonation ASR results are shown in Table 4.
-Table 4: Comparisons of ASR results of dodging attack and impersonation attacks on the LFW dataset. * denote the method that only use the impersonation loss to craft the adversarial face examples.
-[…]
-IR152 FaceNet MF
-Baseline 2.2 18.2 99.2 25.6
-Lagrangian 3.9 26.5 100.0 26.0
-Fine-tuning 3.4 23.2 99.9 25.8
-RZ 0.6 4.4 94.8 15.1
-Pruning 3.4 24.8 100.0 25.4
-PPR 6.8 35.5 100.0 25.9
-Table 4 demonstrates that our proposed pruning method for adversarial examples achieves a significantly smaller decrease in ASR than RZ after pruning 20% of adversarial perturbations, indicating the effectiveness of our proposed approach.
-After pruning and restoration, adversarial perturbations using our proposed Pruning method, both impersonation and dodging ASR of the crafted adversarial face examples are recovered and higher than the Baseline attack method.
-These results demonstrate the effectiveness of our proposed PPR in improving the dodging performance of adversarial attacks on FR without compromising the impersonation attack performance.
-Figure 3: The dodging ASR in different sparsity ratio.
-4.5 Hyperparameter Sensitivity Study
-The sparsity ratio quantifies the proportion of adversarial perturbations that are allowed to be discarded during the pruning stage.
-This ratio greatly impacts the performance of our proposed attack method.
-Hence, we conducted a sensitivity study on the sparsity ratio to analyze its effect on performance of the algorithm.
-We selected the DI-based Lagrangian attack method as the Baseline for pre-training and performed a hyperparameter sensitivity study on LFW with FaceNet as the surrogate model.
-We adjusted the value of $\tilde{\lambda}$ to ensure that the average black-box impersonation ASR results were within a 0.8% absolute difference compared to the corresponding Baseline.
-The dodging ASR results are illustrated in Figure 3.
-Figure 3 illustrates that the dodging ASR of our proposed method initially increases and then decreases as the sparsity ratio increases.
-It is not hard to see the reason; When the sparsity ratio increases, a greater number of adversarial perturbations are pruned, creating more empty regions for the adversarial perturbations that favor dodging attacks in the restoration stage.
-If the sparsity ratio is set to a too high value, an excessive number of adversarial perturbations are allowed to be freed up, resulting in a degradation of performance for the pre-trained adversarial examples.
-Consequently, the performance of adversarial face examples will decrease.
-5 Conclusion
-In this paper, we delve into the issue of multi-identity samples present among adversarial face examples.
-Our research reveals that traditional impersonation attacks often involve multi-identity samples and the success of an traditional impersonation attack may not necessarily imply success in dodging attacks in the black-box setting.
-In order to improve dodging performance without compromising impersonation performance, we proposed a novel attack known as PPR.
-PPR involves pruning the adversarial examples to release some of the adversarial perturbations while tending to maintain the performance of the adversarial examples.
-We then restore the adversarial perturbation in the pruned area, which favors dodging attack, to enhance the performance of dodging attack without harming the performance of the impersonation attack.
-Our extensive experiments demonstrate the effectiveness of our proposed attack method.
-6 Potential Broader Impact
-This paper showcases research dedicated to advancing the field of Machine Learning.
-The proposed attack method could jeopardize the security of FR systems.
-Our aim is to raise awareness through our proposed attack method and reinforce the robustness of FR systems.
-Appendix A Multi-identity Samples among the Benign Face Images
-Multi-identity samples are intriguing samples that can be classifies as multiple classes in FR.
-In this subsection, we will explore the existence of multi-identity samples among the benign face images (i.e.
-the face images without being added adversarial perturbations).
-We randomly select negative face pairs from the entire LFW dataset.
-Subsequently, we use MF as our FR model to extract the embeddings of the face images in each face pair and calculate the cosine similarity between the two images.
-If the cosine similarity surpasses the threshold, both images in the pair are classified as belonging to multiple identities, indicating that they are multi-identity samples.
-Our findings demonstrate the presence of multi-identity samples among the benign face images, as illustrated in Figure 4.
-Figure 4 illustrates that the multi-identity samples among the benign face images closely resemble the appearances of the identities they are classified into, suggesting they are hard samples(Zhou et al., 2023b)(Huang et al., 2020).
-To analyze the cause of this phenomenon, we need to consider the properties of both the multi-identity samples and the FR model.
-Commonly-used FR models are well-trained and capable of correctly classifying the majority of benign face images.
-However, there are some benign face images that the FR model fails to classify accurately, and these samples are referred to as hard samples.
-Multi-identity samples are a specific type of hard sample known as hard negative samples.
-Typically, hard negative samples exhibit a similar appearance(Zeng et al., 2020), indicating that the multi-identity samples among the benign face images share a resemblance.
-Figure 4: The Illustration of the multi-identity samples.
-The images in the first row and third row are the images of different identities.
-The image in the second row are the images of the multi-identities samples that will be recognized as the identities both top and below it by the FR model.
-Appendix B Universality of Multi-identity Samples among Adversarial Face Examples
-The traditional impersonation attacks craft adversarial face example only use the impersonation loss (e.g.
-Equation 5).
-In this subsection, we will investigate the ratio of multi-identity samples and evaluate the effectiveness of traditional impersonation attacks in terms of dodging attacks.
-We utilize the multi-identity sample ratio (MSR) to gauge the proportion of multi-identity samples that can be recognized as both the attacker and victim identities in the adversarial face examples capable of executing successful impersonation attacks.
-MSR be calculate as:
+Let us revisit an early hypothesis posited by Goodfellow et al.
+[14], suggesting that the linear nature of modern DNNs, resembling linear models trained on the same dataset, is the underlying cause of adversarial examples and their surprising transferability [15].
+Based on the hypothesis, the FR model $\mathcal{F}$ can be represented in the following linear form:
+$\mathcal{F}\left(x\right)=\omega x+b$ (16)
+where $\omega$ is the weight vector, and $b$ is the bias term.
+Hence, the absolute variance of the output of the function $\mathcal{F}$ can be articulated as:
+$\begin{split}\rho&=\|\mathcal{F}\left(x+\delta\right)-\mathcal{F}\left(x\right)\|\\ &=\|\omega\left(x+\delta\right)+b-\left(\omega x+b\right)\|\\ &=\|\omega\delta\|\leq\|\omega\|\|\delta\|=\|\omega\|\mathcal{I}\\ \end{split}$ (17)
+The supremum of the absolute variance is given by $\lVert\omega\rVert\mathcal{I}$.
+Within the context of gradient-based adversarial attacks, the parameters of the neural networks are typically held constant.
+Consequently, the vector $\omega$ can be treated as a constant.
+From this, we can deduce that the supremum of the absolute variance in the output of $\mathcal{F}$ is directly proportional to the priority measure $\mathcal{I}$.
+It follows that adversarial perturbations of greater magnitude are more prioritized, as they exert a more pronounced influence on the output of the face recognition model $\mathcal{F}$.
+Appendix D The Calculation Method of the Attack Success Rate
+Due to the different objectives of the impersonation and dodging attacks, the calculation methods for the ASR also vary between the two attack types.
+When it comes to impersonation attacks on FR, the ASR can be computed as:
+${\rm ASR}^{i}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)<t^{i}\right)}{N_{p}}$ (18)
+where $N_{p}$ refers to the total number of face pairs and $t^{i}$ represents the impersonation attack threshold. For dodging attacks on FR systems, the ASR can be computed via the following formula:
+${\rm ASR}^{d}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)>t^{d}\right)}{N_{p}}$ (19)
+where the value of $t^{d}$ represents the threshold for dodging attacks.
+Appendix E The Calculation Method of the Multi-identity Sample Ratio
+Multi-identity Sample Ratio (MSR) is the proportion of multi-identity samples that can be recognized as both the attacker and victim identities in the adversarial face examples capable of executing successful impersonation attacks.
+The detailed calculation method of MSR be expressed as:
 ${\rm MSR}=\frac{|\mathcal{S}_{1}\cap\mathcal{S}_{2}|}{|\mathcal{S}_{2}|}$ (20)
 where $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are the set of the adversarial examples that are identified as the attacker and victim identities, respectively.
-Therefore, the $\mathcal{S}_{2}$ is the set of the adversarial face examples that can perform successful impersonation attack.
-Equation 3 can be rewritten as:
-${\rm ASR}^{i}=\frac{|\mathcal{S}_{2}|}{N_{p}}$ (19)
-We evaluate the MSR, impersonation attack ASR and dodging attack ASR using MF as the surrogate model on the LFW dataset. The results are shown in Table 5.
-Table 5: The results (%) on the LFW dataset. The models on the first row are the victim models.
+Therefore, the $\mathcal{S}_{2}$ is the set of adversarial face examples that can perform a successful impersonation attack.
+Table 4: Comparisons of ASR results (%) of dodging attack and impersonation attacks on the LFW dataset using the negative cosine similarity loss. The models in the second row are the victim models.
+ASRd ASRi
+IR152 FaceNet MF
+Baseline 2.2 16.1 99.4 25.4
+Baseline + Lagrangian 4.0 26.5 100.0 26.0
+Baseline + ours 8.1 39.7 100.0 26.2
+Appendix F Results on Negative Cosine Similarity Loss
+Apart from the normalized $L_{2}$ distance loss (i.e.
+Eq.
+1 and Eq.
+2), the negative cosine similarity loss is also a commonly used loss function in adversarial attacks on FR systems.
+In this section, we will assess the effectiveness of our proposed method regarding the negative cosine similarity loss.
+For impersonation attacks, the negative cosine similarity loss can be expressed as:
+$\mathcal{L}^{i}_{c}=1-\cos\left(\mathcal{F}\left(x^{adv}\right),\mathcal{F}\left(x^{t}\right)\right)$ (21)
+where $\cos\left(e,e^{\prime}\right)$ is the function to calculate the cosine similarity between $e$ and $e^{\prime}$.
+Similarly, the negative cosine similarity loss for dodging attacks can be formulated as follows:
+$\mathcal{L}^{d}_{c}=1+\cos\left(\mathcal{F}\left(x^{adv}\right),\mathcal{F}\left(x^{s}\right)\right)$ (22)
+To evaluate the effectiveness of our proposed method on the negative cosine similarity loss, we use DI as the Baseline to craft adversarial face examples using the negative cosine similarity loss $\mathcal{L}^{i}_{c}$ and $\mathcal{L}^{d}_{c}$ as $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, respectively.
+The performance is demonstrated in Table 4.
+Table 4 demonstrates that the performance of both impersonation and dodging improves after integrating our proposed attack method into the Baseline attack.
+This further highlights the effectiveness of our method regarding the negative cosine similarity loss.
+Table 5: The results (%) on the LFW dataset using MF as the surrogate model. The models on the first row are the victim models.
+ASRd ASRi
+IR152 FaceNet MF
+SA 99.4 45.9 2.1 44.5
+SA + Lagrangian 100.0 49.2 3.4 44.8
+SA + ours 100.0 54.9 4.7 46.4
+Appendix G Results on Sibling-Attack
+Sibling-Attack [29] (abbreviated as SA) is a cutting-edge adversarial attack on FR that demonstrates considerable effectiveness in enhancing the transferability of adversarial attacks on FR.
+SA employs a surrogate model with a two-branch multi-task neural network, which differs from the surrogate models discussed in Section 4.2.
+Consequently, we exclusively consider SA for comparison with our method.
+We aligned the attack setting with the official SA implementation and integrated our proposed Adv-Pruning with SA.
+The FFHQ dataset, provided by the official SA page [^1] https://github.com/Tencent/TFace/tree/Adv-Attack-Defense/security/tasks/Adv-Attack-Defense/Sibling-Attack, is chosen as the dataset for generating adversarial face examples.
+The results are depicted in Table 5.
+Table 5 demonstrates that the incorporation of our proposed method results in enhancements in both dodging and impersonation performance.
+This further demonstrates the effectiveness of our proposed attack method in enhancing the dodging attack performance without compromising the impersonation attack performance.
+Table 6: The results (%) on the LFW dataset using MF as the surrogate model. The models on the first row are the victim models.
 Attack Metric IR152 FaceNet MF
 DI MSR 95.6 75.7 0.8
-[…]
-[…]
+ASRi 18.4 32.9 100.0
+ASRd 2.2 18.2 99.2
 VMI MSR 97.0 88.6 0.3
-[…]
-[…]
+ASRi 13.6 20.2 100.0
+ASRd 1.0 8.4 99.7
 SSA MSR 98.6 92.8 1.5
-[…]
-[…]
+ASRi 13.8 19.6 100.0
+ASRd 0.7 5.9 98.5
 DFANet MSR 98.6 96.6 0.4
-[…]
-[…]
+ASRi 7.0 11.9 100.0
+ASRd 0.2 1.5 99.6
 SIA MSR 97.4 83.9 1.6
-[…]
-[…]
+ASRi 15.7 26.7 100.0
+ASRd 1.0 10.6 98.4
 BSR MSR 98.1 96.8 15.1
-[…]
-[…]
+ASRi 5.4 9.5 100.0
+ASRd 0.4 3.7 84.9
 BPFA MSR 96.2 92.1 2.3
-[…]
-[…]
-Table 5 demonstrate that most of the crafted adversarial face examples are multi-identity samples when the victim model are the black-box model.
-This indicates that most adversarial face examples generated through traditional impersonation attacks are unable to attain a successful dodging attack in the black-box setting.
-Nevertheless, in the white-box setting, the majority of adversarial face examples capable of executing successful impersonation attacks also demonstrate success in dodging attacks.
-To analyze the reason, we need to consider the metric used to determine whether two face images belong to the same identity in FR.
-FR is an open-set task, so we utilize the distance in the embedding space to make this determination.
-Figure 5 illustrates two decision boundaries for each FR model, one for attacker identity and one for victim identity.
-In the white-box setting, if we generate adversarial examples using Equation 5, these examples can enter a space in which they are identified as the victim identity but not as the attacker identity.
-However, the decision boundary of the black-box model differs from that of the surrogate model.
-In the black-box setting, most adversarial examples are found between the decision boundaries of the black-box model, resulting in the majority of adversarial examples crafted using Equation 5 being multi-identity samples.
-This demonstrates that the adversarial face examples are positioned near the decision boundary, aligning with findings from prior research of adversarial attacks on image classification(Cao & Gong, 2017).
-Figure 5: Illustration of the decision boundaries of FR models.
-The star in the left and right are the initial image of the adversarial face examples (i.e.
-the attacker image) and the final adversarial example, respectively.
-The solid and dashed lines are the decision boundaries for the surrogate and black-box models, respectively.
-For both the surrogate and black-box models, the lines in the left and right are the decision boundary of the victim and attacker, respectively.
+ASRi 15.8 17.8 100
+ASRd 0.9 4.6 97.7
+Appendix H Results on MSR of Previous Impersonation Attacks
+Following the same experimental setting of Section 4.4, we conduct experiments on various previous impersonation attacks. The results are shown in Table 6.
+As indicated by Table 6, the majority of the crafted adversarial face examples are multi-identity samples in the black-box setting.
+This observation suggests that the majority of adversarial face examples, which are generated through the previous impersonation attacks, fail to achieve a successful dodging attack in the black-box setting.

### v4

@@
 Rethinking Impersonation and Dodging Attacks on Face Recognition Systems
-Abstract
+Abstract.
 Face Recognition (FR) systems can be easily deceived by adversarial examples that manipulate benign face images through imperceptible perturbations.
 Adversarial attacks on FR encompass two types: impersonation (targeted) attacks and dodging (untargeted) attacks.
-Previous methods often achieve a successful impersonation attack on FR; However, it does not necessarily guarantee a successful dodging attack on FR in the black-box setting.
+Previous methods often achieve a successful impersonation attack on FR, however, it does not necessarily guarantee a successful dodging attack on FR in the black-box setting.
 In this paper, our key insight is that the generation of adversarial examples should perform both impersonation and dodging attacks simultaneously.
@@
 As a result, we can maintain the impersonation capabilities of original adversarial examples while effectively enhancing dodging capabilities.
-Comprehensive experiments demonstrate the superiority of our method compared with state-of-the-art adversarial attacks.
-1 Introduction
-Figure 1: Top: previous methods that achieve a successful impersonation attack on FR cannot guarantee a successful dodging attack on FR in the black-box setting.
+Comprehensive experiments demonstrate the superiority of our method compared with state-of-the-art adversarial attack methods.
+Face Recognition, Adversarial Attacks, Adversarial Attacks on Face Recognition, Impersonation Attacks, Dodging Attacks
+copyright: acmlicensedjournalyear: 2024doi: 10.1145/3664647.3681440conference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10submissionid: 430ccs: Computing methodologies Biometrics
+1. Introduction
+Figure 1.
+Top: previous methods that achieve a successful impersonation attack on FR cannot guarantee a successful dodging attack on FR in the black-box setting.
 In contrast, we present Adv-Pruning, including Priming, Pruning, and Restoration Stages, to perform both impersonation and dodging attacks simultaneously.
-Bottom (left): natural Multi-identity Samples (MS).
+Bottom (left): benign Multi-identity Samples (MS).
 Bottom (right): the dodging Attack Success Rate (%) between the previous methods and Adv-Pruning on multiple models.
-Thanks to the ceaseless advancements in deep learning, Face Recognition (FR) has achieved exceptional performance [41; 49; 9; 1; 2; 26].
-However, the vulnerability of existing FR models to adversarial attacks poses a significant threat to their security.
+Thanks to the ceaseless advancements in deep learning, Face Recognition (FR) has achieved exceptional performance (Schroff et al., 2015; Wang et al., 2018; Deng et al., 2022; An et al., 2021; Boutros et al., 2023; Li et al., 2023a).
+However, the vulnerability of existing FR models poses a significant threat to their security (Das et al., 2021; Zhou et al., 2022a, b, 2023c; Narayan et al., 2023; Wang et al., 2024b; Zhou et al., 2024b), with adversarial attacks being one of the key concerns.
 Hence, there is an urgent need to enhance the performance of adversarial face examples to expose more blind spots in FR models.
 As a result, several research endeavors have been directed towards this realm.
-A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness [40; 62; 7; 19; 42], transferability [70; 29; 72; 71], and physical attack capability [64; 63; 28].
+A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness (Qiu et al., 2020; Yang et al., 2021; Cherepanova et al., 2021; Hu et al., 2022; Shamshad et al., 2023), transferability (Zhong and Deng, 2021; Li et al., 2023d; Zhou et al., 2023b, 2024a), and physical attack capability (Yin et al., 2021; Yang et al., 2023; Li et al., 2023c).
 These efforts contribute to enhancing the effectiveness of adversarial attacks on FR.
@@
 In the realm of adversarial attacks on image classification, a successful impersonation attack typically implies a successful dodging attack.
-However, FR is an open-set task [49; 9], which is quite different from image classification.
+However, FR is an open-set task (Wang et al., 2018; Deng et al., 2022), which is quite different from image classification.
 In the real-world deployment of FR systems, accurately predicting the class probability of identities presents an extreme challenge.
@@
 If the distance falls below a predefined threshold, the two images are recognized as belonging to the same identity; otherwise, they are classified as different identities.
-Based on the measurement of FR, there are two decision boundaries for each FR model when crafting adversarial examples as shown in Fig.
-1.
-As a result, there exists the natural samples that can be classified as two different identities in theory.
-We denote these samples as multi-identity samples (refer to Section 4.4).
+Based on the measurement of FR, there are two decision boundaries for each FR model when crafting adversarial examples as shown in Figure 1.
+As a result, there exists the benign samples that can be classified as two different identities in theory.
+We denote these samples as multi-identity samples.
 The existence of multi-identity samples implies that a successful impersonation attack on FR does not necessarily guarantee a successful dodging attack on FR.
-Existing research indicates that an adversarial sample was located near the decision boundary [18; 4].
+Existing research indicates that an adversarial sample was located near the decision boundary (Heo et al., 2019; Cao and Gong, 2017).
 Suppose we generate adversarial face examples using previous methods.
-In the white-box setting, both the structures and parameters of the victim models are known, enabling the generation of adversarial face examples that can cross the decision boundaries of both attacker and victim, as shown in Fig.
-1.
+In the white-box setting, both the structures and parameters of the victim models are known, enabling the generation of adversarial face examples that can cross the decision boundaries of both attacker and victim, as shown in Figure 1.
 However, in the black-box setting, the decision boundaries of black-box models differ from those of the surrogate models.
@@
 As such, the majority of adversarial face examples crafted by previous methods, which can successfully perform impersonation attacks, fail to perform dodging attacks in the black-box setting.
-Figure 2: Overview of our Adv-Pruning attack framework, which consists of Priming, Pruning, and Restoration stages.
-(a) During the Priming stage, we optimize the adversarial examples to ensure they have sufficient attack performance.
-(b) In the Pruning stage, we propose Adversarial Priority Quantification to quantify the priority of adversarial perturbations.
-Subsequently, we sparsify the adversarial perturbations based on the quantified priorities.
-(c) In the Restoration stage, we present Biased Gradient Adaptation to introduces gradient perturbations biased to dodging attacks on the sparsified regions.
 In this paper, we propose a novel attack method, termed as Adversarial Pruning (Adv-Pruning).
@@
 In the Priming stage, we optimize the adversarial examples to ensure adequate attack potential.
-In the Pruning stage, with considering the pruning concept in model compression, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations using an priority measure which is directly proportional to the supremum of the absolute model output variances.
+In the Pruning stage, considering the pruning concept in model compression, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations using a priority measure that is directly proportional to the supremum of the absolute model output variances.
 After processing by Adversarial Priority Quantification, we prune the adversarial face examples to free up less prioritized adversarial perturbations.
 In the Restoration stage, we propose Biased Gradient Adaptation to add biased gradient perturbations favoring dodging attacks on the pruned regions to adapt the adversarial face examples into the space that can be classified as the victim while remaining unidentifiable as the attacker, thereby enhancing the dodging performance of the adversarial face examples without compromising the prioritized features of original adversarial perturbations.
-As illustrated in the top of Fig.
-1, after undergoing these stages, the adversarial face example generated by our proposed method can successfully traverse the decision boundaries of both the attacker and victim of the black-box model, achieving successful black-box impersonation and dodging attacks.
+As illustrated in the top of Figure 1, after undergoing these stages, adversarial face examples generated by our method can successfully traverse the decision boundaries of both the attacker and victim of the black-box model, achieving successful black-box impersonation and dodging attacks.
 Our main contributions are summarized as follows:
@@
 Moreover, our presented method could be plugged into various FR systems and adversarial attack methods.
-2 Related Work
-2.1 Adversarial Attacks
-The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images to deceive machine learning systems and cause them to make mistakes [47; 14].
+2. Related Work
+2.1. Adversarial Attacks
+The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images to deceive machine learning systems and cause them to make mistakes (Szegedy et al., 2014; Goodfellow et al., 2015).
 The existence of adversarial examples poses a significant threat to the security of current machine learning systems.
-Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems [33; 69; 30; 55; 34; 43; 12; 73; 68; 37].
-To improve the performance of black-box adversarial attacks, DI [60] applies random transformations to adversarial examples in each iteration to achieve a data augmentation effect.
-VMI-FGSM [53] employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance.
-SSA [33] transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them.
-SIA [54] applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculation.
-BSR [51] divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation.
-DA [16] utilizes dispersion amplification to enhance the multi-task attack capability of adversarial attacks.
-Despite their gratifying progress, these studies neglect the consideration of pruning adversarial examples through introducing pruning methods into the realm of adversarial attacks.
+Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems (Long et al., 2022; Liang et al., 2023; Lu et al., 2023; Shayegani et al., 2024; Ge et al., 2023a; Zhou et al., 2023a; Zhang et al., 2022b; Mingxing et al., 2021).
+To improve the performance of black-box adversarial attacks, DI (Xie et al., 2019) applies random transformations to adversarial examples in each iteration to achieve a data augmentation effect.
+VMI-FGSM (Wang and He, 2021) employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance.
+SSA (Long et al., 2022) transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them.
+SIA (Wang et al., 2023) applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculation.
+BSR (Wang et al., 2024a) divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation.
+DA (Haleta et al., 2021) utilizes dispersion amplification to enhance the multi-task attack capability of adversarial attacks.
+Despite their gratifying progress, these studies neglect the consideration of pruning adversarial examples by introducing pruning methods into the realm of adversarial attacks.
 In our research, we propose a novel pruning method capable of identifying and freeing up the adversarial perturbations with minimal impact on absolute model output variances, thereby sparsifying regions for adding adversarial perturbations with the aim of dodging capabilities improvement.
-2.2 Adversarial Attacks on Face Recognition
-Based on the restriction of the adversarial perturbations, adversarial attacks on FR can be classified into two categories: restricted attacks [32; 61; 10; 74; 35; 31; 5] and unrestricted attacks [3; 6; 65; 8; 44; 46; 48; 56; 57].
+2.2. Adversarial Attacks on Face Recognition
+Based on the restriction of the adversarial perturbations, adversarial attacks on FR can be classified into two categories: restricted attacks and unrestricted attacks.
 Restricted attacks on FR are the attacks that generate adversarial examples in a restricted bound (e.g.
 $L_{p}$ bound).
-To enhance the transferability of adversarial attacks on FR, [70] propose DFANet, which applies dropout on the feature maps of the convolutional layers to achieve ensemble-like effects.
-In addition, [72] introduces BPFA, which further improves the transferability of adversarial attacks on FR by incorporating beneficial perturbations [58] on the feature maps of the FR models, resulting in hard model augmentation effects.
-[29] leverages extra information from FR-correlated tasks and uses a multi-task optimization framework to enhance the transferability of crafted adversarial examples.
+To enhance the transferability of adversarial attacks on FR, Zhong and Deng (Zhong and Deng, 2021) propose DFANet, which applies dropout on the feature maps of the convolutional layers to achieve ensemble-like effects.
+In addition, Zhou et al.
+(Zhou et al., 2023b) introduces BPFA, which further improves the transferability of adversarial attacks on FR by incorporating beneficial perturbations (Wen and Itti, 2020) on the feature maps of the FR models, resulting in hard model augmentation effects.
+Li et al.
+(Li et al., 2023d) leverages extra information from FR-related tasks and uses a multi-task optimization framework to enhance the transferability of crafted adversarial examples.
+Zhou el al.
+(Zhou et al., 2024c) propose an adversarial attack that can attack both FR and Face Anti-spoofing (FAS) models simultaneously, aiming to enhance the practicability of adversarial attacks on FR systems.
 The unrestricted adversarial attacks on FR are the attacks that generate adversarial examples without the restriction of a predefined perturbation bound.
-They mainly focus on physical attacks [59; 63; 28], attribute editing [40; 23] and generating adversarial examples based on makeup transfer [64; 19; 42].
+They mainly focus on physical attacks (Xiao et al., 2021; Yang et al., 2023; Li et al., 2023c), attribute editing (Qiu et al., 2020; Jia et al., 2022) and generating adversarial examples based on makeup transfer (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023).
 The existing literature on both restricted and unrestricted adversarial attacks on FR systems has successfully enhanced the performance of these attacks.
-Nevertheless, it remains under-explored in the correlation between impersonation and dodging attacks.
-This paper elegantly addresses this by investigating the correlation between impersonation and dodging attacks and introducing a novel attack method that bolsters the dodging capabilities while preserving the impersonation capabilities of previous methods.
-3 Methodology
-3.1 Problem Formulation
+Nevertheless, it remains under-explored in the correlation between impersonation and dodging attacks on FR.
+This paper addresses this by investigating the correlation between impersonation and dodging attacks and introducing a novel attack method that bolsters the dodging capabilities while preserving the impersonation capabilities of previous methods.
+3. Methodology
+3.1. Problem Formulation
 Let $\mathcal{F}^{vct}(x)$ denote the FR model used by the victim to extract the embedding from a face image $x$.
@@
 By contrast, the objective of the dodging attacks proposed in this study is to render $\mathcal{F}^{vct}(x)$ unable to identify $x^{adv}$ as $x^{s}$, while simultaneously ensuring that $x^{adv}$ bears a visual resemblance to $x^{s}$.
-For the sake of clarity and conciseness, the detailed optimization objectives for both impersonation and dodging attacks are provided in the appendices.
+For the sake of clarity and conciseness, the detailed optimization objectives for both impersonation and dodging attacks are provided in the supplementary.
 Few works explore the correlation between impersonation and dodging attacks on FR.
 In the following, we delve into the correlation between these two types of attacks and propose a novel method to enhance dodging attacks while maintaining impersonation attacks.
-An overview of the proposed method is illustrated in Fig.
-2.
-As depicted in Fig.
-2, our proposed method is structured into three stages: Priming, Pruning, and Restoration.
+An overview of the proposed method is illustrated in Figure 2.
+As depicted in Figure 2, our proposed method is structured into three stages: Priming, Pruning, and Restoration.
 Through the sequential application of these stages, we are able to generate adversarial examples that exhibit a potent combination of impersonation and dodging attack capabilities.
-3.2 Exploring the Impersonation and Dodging Attack on Face Recognition
+Figure 2.
+Overview of our Adv-Pruning attack framework, which consists of Priming, Pruning, and Restoration stages.
+(a) During the Priming stage, we optimize the adversarial examples to ensure they have sufficient attack performance.
+(b) In the Pruning stage, we propose Adversarial Priority Quantification to quantify the priority of adversarial perturbations.
+Subsequently, we sparsify the adversarial perturbations based on the quantified priorities.
+(c) In the Restoration stage, we present Biased Gradient Adaptation to introduce gradient perturbations biased to dodging attacks on the sparsified regions.
+3.2. Exploring the Impersonation and Dodging Attack on Face Recognition
 In most cases, the victim model $\mathcal{F}^{vct}$ is not accessible to the attacker, making it extremely challenging to optimize the objectives for black-box attacks directly.
-To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversarial examples that can be transferred to the victim model for an effective attack [66; 69; 39; 54; 51; 38; 11; 13; 27; 5].
+To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversarial examples that can be transferred to the victim model for an effective attack (Yuan et al., 2022; Zhang et al., 2022a; Naseer et al., 2023; Wang et al., 2023, 2024a; kanth Nakka and Salzmann, 2021; Gao et al., 2023; Ge et al., 2023b; Li et al., 2023b; Chen et al., 2023).
 For impersonation attacks, the loss can be formulated as follows:
-$\mathcal{L}^{i}=\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{t}\right)\right)\|^{2}_{2}$ (1)
+(1) $\mathcal{L}^{i}=\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{t}\right)\right)\|^{2}_{2}$
 where $\phi(x)$ represents the operation that normalizes $x$.
@@
 The loss function of dodging attacks can be formulated as:
-$\mathcal{L}^{d}=-\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{s}\right)\right)\|^{2}_{2}$ (2)
+(2) $\mathcal{L}^{d}=-\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{s}\right)\right)\|^{2}_{2}$
 As the FR task is an open-set task, it is impractical to predict the classes of users during the practical deployment of the FR model.
@@
 Based on the identification method in FR, multi-identity samples exist theoretically.
-Our experiments verify the existence of such samples among benign face images.
+Our experiments verify the existence of such samples among benign face images (see the supplementary).
 The existence of multi-identity samples raises a question:
@@
 To this end, we generate adversarial face examples using the previous impersonation attack and evaluate its dodging Attack Success Rate (ASR).
-Our experiment confirms that the majority of adversarial examples crafted through previous methods, which are successful in performing impersonation attacks, fail to successfully execute dodging attacks in the black-box setting (see Section 4.4).
+Our experiment confirms that the majority of adversarial examples crafted through previous methods, which are successful in performing impersonation attacks, fail to successfully execute dodging attacks in the black-box setting.
 Nonetheless, in real-world adversarial attacks, attackers do not want the adversarial face examples to be recognized as themselves, as this may lead to legal consequences.
@@
 To accomplish this objective, a straightforward approach is to generate adversarial face examples using a multi-task attack strategy.
-In the following, we will take the Lagrangian attack strategy as the example for its simplicity.
+In the following, we will take the Lagrangian attack strategy as an example for its simplicity.
 The Lagrangian attack strategy utilizes the following loss function to craft adversarial examples:
-$\mathcal{L}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$ (3)
+(3) $\mathcal{L}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$
 However, due to the conflict between the optimization between $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, there exists a trade-off between the performance of impersonation and dodging performance, leading to subpar performance (See Section 4.4).
 Suppose we can mitigate the trade-off, we will achieve a better dodging performance while maintaining the impersonation performance.
-3.3 Adversarial Pruning Attack
+3.3. Adversarial Pruning Attack
 To accomplish this objective, a straightforward approach is to fine-tune the adversarial face examples generated by the Lagrangian attack with a lower $\lambda$ value in order to enhance the performance of dodging attacks.
@@
 In the following, we will introduce our proposed Adv-Pruning based on Lagrangian attack in detail.
-In the Priming Stage, we utilize Eq.
-3 as the Priming loss $\mathcal{L}^{p}$ to craft the adversarial face examples:
-$x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)\right)$ (4)
+In the Priming Stage, we utilize Equation (3) as the Priming loss $\mathcal{L}^{p}$ to craft the adversarial face examples:
+(4) $x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)\right)$
 where $t$ is the iteration of the optimization process of adversarial examples, and $\beta$ is the step size when optimizing the adversarial face examples in the Priming stage, and $\prod\left(x\right)$ is the projection function that projects $x$ onto the $L_{p}$ norm bound.
@@
 A lower magnitude implies a lesser impact on the performance of the adversarial examples generated after sparsification, as the supremum of absolute model output variances is directly proportional to the magnitude of the adversarial perturbations.
-The proof is in the appendices.
+The proof is in the supplementary.
+Figure 3. The ASR on victim models trained by multiple algorithms with IR152 as the surrogate model.
 Let the adversarial examples be $x^{adv}$. The formula to calculate the priority can be expressed as:
-$\mathcal{I}=|x^{adv}-x^{s}|$ (5)
+(5) $\mathcal{I}=|x^{adv}-x^{s}|$
 where $\mathcal{I}\in\mathbb{R}^{\rm CHW}$. ${\rm C}$, ${\rm H}$, and ${\rm W}$ are the channel number, height, and width of the face images, respectively.
 Once the priority values of the adversarial perturbations are quantified, we employ these values to release less prioritized adversarial perturbations.
-Let $\kappa$ be the sparsity ratio for pruning the adversarial face examples that measure the ratio of perturbations to be set into zero.
+Let $\kappa$ be the sparsity ratio for pruning the adversarial face examples that measure the ratio of perturbations to be set to zero.
 Let $s={\rm CHW}$ be the number of adversarial perturbation elements.
 We arrange the elements in a flattened vector of $\mathcal{I}$ in ascending order (from the lowest to the highest):
-$\mathcal{Q}={\rm Sort}\left(\Psi\left(\mathcal{I}\right)\right)$ (6)
+(6) $\mathcal{Q}={\rm Sort}\left(\Psi\left(\mathcal{I}\right)\right)$
 where $\Psi$ is the flatten operation.
 Let $\mathcal{W}$ be the set of the elements of the adversarial perturbations to be pruned. Given the priority calculation method for pruning, the value of $\mathcal{W}$ can be calculated as follows:
-$\mathcal{W}=\mathcal{Q}\left[:\kappa s\right]$ (7)
+(7) $\mathcal{W}=\mathcal{Q}\left[:\kappa s\right]$
 where the colon denotes the slice operation to obtain the first $\kappa s$ elements. The pruning mask, which has the same shape as $\mathcal{I}$, can be obtained by utilizing $\mathcal{W}$:
-$\mathcal{M}_{i,j,k}=\begin{cases}0,&\text{if }\mathcal{I}_{i,j,k}\in\mathcal{W}\\ 1,&\text{if }\mathcal{I}_{i,j,k}\notin\mathcal{W}\end{cases}$ (8)
+(8) $\mathcal{M}_{i,j,k}=\begin{cases}0,&\text{if }\mathcal{I}_{i,j,k}\in\mathcal{W}\\ 1,&\text{if }\mathcal{I}_{i,j,k}\notin\mathcal{W}\end{cases}$
 By utilizing the mask, we can apply the following formula to prune the adversarial example:
-$\bar{x}^{adv}=x^{s}+\left(x^{adv}-x^{s}\right)\odot\mathcal{M}$ (9)
+(9) $\bar{x}^{adv}=x^{s}+\left(x^{adv}-x^{s}\right)\odot\mathcal{M}$
 where $\bar{x}^{adv}$ is the adversarial face example after pruning.
+Figure 4. Comparisons of ASR (%) on LFW with adversarial robust models as victim models.
 Biased Gradient Adaptation.
 During the Restoration stage, we restore the adversarial face examples in the previously pruned region using our proposed Biased Gradient Adaptation.
-Biased Gradient Adaptation using the following loss function to craft gradient biased to the dodging attacks to adapt the crafted adversarial examples into the space that favors dodging attacks.
-$\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$ (10)
+Biased Gradient Adaptation uses the following loss function to craft gradient biased to the dodging attacks to adapt the crafted adversarial examples into the space that favors dodging attacks.
+(10) $\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$
 where $\tilde{\lambda}$ is a weight that is lower than $\lambda$ that is objective for crafting adversarial face examples that favor dodging attacks.
 The mask representing the regions for restoring the adversarial examples can be denoted as:
-$\mathcal{A}=1-\mathcal{M}$ (11)
+(11) $\mathcal{A}=1-\mathcal{M}$
 Subsequently, we utilize the following formula to restore the pruned adversarial face examples:
-$\begin{split}x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{n}+\mathcal{A}\odot\left(x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)-\bar{x}^{adv}\right)\right)\end{split}$ (12)
+(12) $\begin{split}x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{n}+\mathcal{A}\odot\left(x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)-\bar{x}^{adv}\right)\right)\end{split}$
 where $\gamma$ is the step size when optimizing the adversarial face examples in the Restoration stage, $\bar{x}^{adv}$ is the adversarial example crafted by the Priming Stage, and $\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}$ is the biased gradient.
-The pseudo-code of our proposed method based on the Lagrangian attack is illustrated in the appendices.
-4 Experiments
-4.1 Experimental Setting
+The pseudo-code of our proposed method based on the Lagrangian attack is illustrated in the supplementary.
+4. Experiments
+4.1. Experimental Setting
 Datasets.
-We opt to use the LFW [20], CelebA-HQ [24], and FFHQ [25] datasets for our experiments.
+Face images play a pivotal role in multimedia processing applications.
+Therefore, the research on adversarial attacks on FR has a significant impact on security and privacy in multimedia processing.
+We opt to use the LFW (Huang et al., 2007), CelebA-HQ (Karras et al., 2018), FFHQ (Karras et al., 2019), and BUPT-Balancedface (Wang et al., 2019) datasets for our experiments.
 LFW serves as an unconstrained face dataset for FR.
-CelebA-HQ and FFHQ consists of high-quality images.
-The LFW and CelebA-HQ utilized in our experiments are identical to those employed in [72; 71], while FFHQ is the corresponding dataset provided by the Sibling-Attack official page, ensuring the consistency for analysis.
-Figure 3: The ASR on FR models trained by multiple algorithms.
-Figure 4: Comparisons of ASR (%) on LFW with adversarial robust models as victim models.
-Table 1: Comparisons of dodging ASR (%) results for attacks on the LFW and CelebA-HQ datasets.
+CelebA-HQ and FFHQ consist of high-quality images.
+The BUPT-BalancedFace dataset is designed to address the racial imbalance present in existing face datasets.
+The LFW and CelebA-HQ utilized in our experiments are identical to those employed in (Zhou et al., 2023b, 2024a), while FFHQ is the corresponding dataset provided by the Sibling-Attack official page, ensuring the consistency for analysis.
+For BUPT-Balancedface, we randomly selected 1,000 pairs from the dataset for evaluating the performance of the adversarial examples.
+Table 1.
+Comparisons of dodging ASR (%) results for attacks on the LFW and CelebA-HQ datasets.
 The surrogate models are presented in the first column, and the victim models are listed in the second row.
-The numbers before and after the slash represent the results of the baseline attack and the attack that combines the baseline attack with our proposed attack, respectively.
+The numbers before and after the slash represent the results of the baseline and the attack that combines the baseline with our proposed method, respectively.
 LFW CelebA-HQ
 Surrogate Model Attack IR152 FaceNet MF IR152 FaceNet MF
-IR152 [17] DI 95.4 / 100.0 5.8 / 11.3 0.2 / 0.8 87.9 / 100.0 8.4 / 16.4 0.4 / 1.2
+IR152 (He et al., 2016) DI 95.4 / 100.0 5.8 / 11.3 0.2 / 0.8 87.9 / 100.0 8.4 / 16.4 0.4 / 1.2
 VMI 92.7 / 100.0 17.3 / 32.5 1.2 / 9.5 92.2 / 100.0 14.3 / 27.9 1.2 / 4.1
@@
 BPFA 92.6 / 100.0 1.7 / 7.3 0.0 / 1.2 90.4 / 100.0 2.1 / 8.1 0.1 / 0.8
-FaceNet [41] DI 5.3 / 10.3 99.8 / 99.9 3.1 / 10.3 1.5 / 3.1 99.4 / 99.9 1.8 / 4.7
+FaceNet (Schroff et al., 2015) DI 5.3 / 10.3 99.8 / 99.9 3.1 / 10.3 1.5 / 3.1 99.4 / 99.9 1.8 / 4.7
 VMI 9.7 / 14.3 99.8 / 99.9 6.2 / 13.2 3.1 / 7.1 99.3 / 99.8 3.6 / 9.3
@@
 BPFA 4.7 / 16.8 98.6 / 100.0 1.6 / 15.0 1.1 / 4.2 99.0 / 100.0 0.6 / 5.1
-MF [9] DI 2.2 / 7.3 18.2 / 36.4 99.2 / 100.0 0.1 / 2.5 12.1 / 31.3 95.2 / 100.0
+MF (Deng et al., 2022) DI 2.2 / 7.3 18.2 / 36.4 99.2 / 100.0 0.1 / 2.5 12.1 / 31.3 95.2 / 100.0
 VMI 1.0 / 2.8 8.4 / 20.9 99.7 / 100.0 0.2 / 0.4 5.2 / 15.0 98.2 / 100.0
@@
 Face Recognition Models.
-The normal trained FR models employed in our experiments include IR152 [17], FaceNet [41], MobileFace (abbreviated as MF) [9], ArcFace [9], CircleLoss [45], CurricularFace [22], MagFace [36], MV-Softmax [52], and NPCFace [67].
-IR152, FaceNet, and MF are identical to those used in [64; 19; 72; 71].
-ArcFace, CircleLoss, CurricularFace, MagFace, MV-Softmax, and NPCFace are the official models available in FaceX-ZOO [50].
-Additionally, we incorporate adversarial robust FR models in our experiments, denoted as IR152adv, FaceNetadv, and MFadv, which are identical to those used in [72].
+The normal trained FR models employed in our experiments include IR152 (He et al., 2016), FaceNet (Schroff et al., 2015), MobileFace (abbreviated as MF) (Deng et al., 2022), ArcFace (Deng et al., 2022), CircleLoss (Sun et al., 2020), CurricularFace (Huang et al., 2020), MagFace (Meng et al., 2021), MV-Softmax (Wang et al., 2020), and NPCFace (Zeng et al., 2020).
+IR152, FaceNet, and MF are identical to those used in (Yin et al., 2021; Hu et al., 2022; Zhou et al., 2023b, 2024a).
+ArcFace, CircleLoss, CurricularFace, MagFace, MV-Softmax, and NPCFace are the official models available in FaceX-ZOO (Wang et al., 2021).
+Additionally, we incorporate adversarial robust FR models in our experiments, denoted as IR152adv, FaceNetadv, and MFadv, which are identical to those used in (Zhou et al., 2023b).
 For calculating the ASR in impersonation and dodging attacks, we choose the thresholds based on FAR@0.001 on the entire LFW dataset.
@@
 We use ASRi and ASRd to denote impersonation and dodging ASR, respectively.
-The detailed calculation methods for ASRi and ASRd are provided in the appendices.
+The detailed calculation methods for ASRi and ASRd are provided in the supplementary.
 Compared methods.
 Our proposed attack is a restricted attack method that aims to maliciously attack FR systems to expose more blind spots of them.
-It is not fair to compare our proposed method with unrestricted attacks that do not limit the magnitude of the adversarial perturbations.
-Therefore, we choose restricted attacks on FR that aim to maliciously attack FR systems [70] [72] [29] and state-of-the-art transfer attacks [60] [33] [54] [51] as our baseline.
-Table 2: Comparisons of ASR (%) with multi-task attacks on LFW dataset. Models in the second row are victim models.
+It is not fair to compare our method with unrestricted attacks that do not limit the magnitude of the adversarial perturbations.
+Therefore, we choose restricted attacks on FR that aim to maliciously attack FR systems (Zhong and Deng, 2021) (Zhou et al., 2023b) (Li et al., 2023d) (Yang et al., 2020) and state-of-the-art transfer attacks (Xie et al., 2019) (Long et al., 2022) (Wang et al., 2023) (Wang et al., 2024a) as our baseline.
+Table 2. Comparisons of ASR (%) with multi-task attacks on LFW dataset. Models in the second row are victim models.
 ASRd ASRi
@@
 DA + Ours 17.5 44.9 99.4 37.8
-4.2 Comparison Study
+4.2. Comparison Study
 We compare our proposed attack method with the state-of-the-art attacks on multiple FR models and datasets.
-Several adversarial examples are illustrated in Fig.
-5.
+Several adversarial examples are illustrated in Figure 5.
 The attack performance results are shown in Table 1.
@@
 Table 2 underscores the effectiveness of our method in enhancing the dodging performance of multi-task attacks while maintaining the impersonation performance.
-To further validate our proposed attack method on additional FR models, we selected SIA [54] as Baseline and IR152 as the surrogate model.
+To further validate our proposed attack method on additional FR models, we selected SIA (Wang et al., 2023) as Baseline and IR152 as the surrogate model.
 The experimental settings are consistent with those described in Table 1.
-The dodging ASR across multiple FR models is demonstrated in Fig.
-3.
-As depicted in Fig.
-3, the dodging ASR improves on multiple FR models after integrating our proposed method, further confirming the effectiveness of our attack.
-Figure 5: The Illustration of adversarial examples crafted by various attacks.
-First column: some attacker images.
+The dodging ASR across multiple FR models is demonstrated in Figure 3.
+As depicted in Figure 3, the dodging ASR improves on multiple FR models after integrating our proposed method, further confirming the effectiveness of our attack.
+Figure 5.
+The Illustration of adversarial examples crafted by various attacks.
+First column: attacker images.
 Last column: the corresponding victim images.
-The second to fifth columns exhibit the corresponding adversarial face examples crafted by BPFA, BPFA + ours, BSR, and BSR + ours, respectively.
+The second to fifth columns exhibit the corresponding adversarial examples of BPFA, BPFA + Ours, BSR, and BSR + Ours, respectively.
 In practical application scenarios, victims can employ adversarial robust models to defend against adversarial attacks.
@@
 In this study, we generate adversarial examples on the LFW dataset using MF as the surrogate model and assess the performance of various attacks on the adversarial robust models.
-The results are presented in Fig.
-4.
-The letters following the en dash represent the surrogate models, with ’I’, ’F’, and ’M’ corresponding to IR152adv, FaceNetadv, and MFadv, respectively.
-Fig.
-4 illustrates that the inclusion of our proposed method leads to improvements in both dodging and impersonation performance.
+The results are presented in Figure 4.
+The letters following the en dash represent the surrogate models, with ‘I’, ‘F’, and ‘M’ corresponding to IR152adv, FaceNetadv, and MFadv, respectively.
+Figure 4 illustrates that the inclusion of our proposed method leads to improvements in both dodging and impersonation performance.
 These results serve as evidence of the effectiveness of our proposed method on adversarial robust models.
@@
 To assess the effectiveness of our proposed attack under JPEG compression, we utilize DI as the baseline attack and MF as the surrogate model, evaluating the attack performance on ArcFace and CurricularFace models with experimental settings consistent with those described in Table 1.
-The results are illustrated in Fig.
-6.
+The results are illustrated in Figure 6.
 These results demonstrate that across varying levels of JPEG compression, our proposed attack method consistently outperforms the baseline attack, thereby highlighting its effectiveness under JPEG compression.
-The experimental results on negative cosine similarity loss, and Sibling-Attack are presented in the appendices.
-4.3 Ablation Study
+The experimental results on negative cosine similarity loss, Sibling-Attack, and LGC are presented in the supplementary.
+4.3. Ablation Study
+Table 3. Comparisons of ASR (%) results of dodging and impersonation attacks on the LFW dataset. The models in the second row are the victim models.
+ASRd ASRi
+Attack IR152 FaceNet MF
+Baseline 2.2 18.2 99.2 25.6
+Lagrangian 3.9 26.5 100.0 26.0
+Fine-tuning 4.2 26.3 100.0 25.6
+RZ 0.6 4.4 94.8 15.1
+Pruning 3.4 24.8 100.0 25.4
+Adv-Pruning 5.4 32.5 100.0 26.3
 To delve into the properties of our proposed attack method, we conducted an ablation experiment using DI as the Baseline attack, with MF serving as the surrogate model on the LFW dataset.
@@
 We applied this method and our pruning method to free up 20% of the adversarial perturbations crafted by the Lagrangian attack.
-For Fine-tuning, we employed the Lagrangian attack as the method to further optimize the Lagrangian adversarial examples with a lower $\lambda$.
+For Fine-tuning, we utilized the Lagrangian attack, setting the weight of the impersonation loss to 12.0, as a method to further optimize the Lagrangian adversarial examples that were initially crafted with an impersonation loss weight of 6.0.
 The dodging attack ASR and average black-box impersonation ASR results are shown in Table 3.
@@
 We conduct a hyperparameter sensitivity study on LFW using FaceNet as the surrogate model, and adjust the value of $\tilde{\lambda}$ to ensure that the average black-box impersonation ASR results were within a 0.4% absolute difference compared to the Baseline.
-The dodging ASR results are illustrated in the right plot of Fig.
-7.
+The dodging ASR results are illustrated in the right plot of Figure 7.
 The results illustrate that the dodging ASR of our proposed method initially increases and then decreases as the sparsity ratio increases.
@@
 Consequently, the performance of adversarial face examples will decrease.
-Figure 6: The dodging ASR under various JPEG Q values.
-4.4 Analytical Study
-Multi-identity Samples among the Natural Face Images: Multi-identity samples are intriguing samples that can be classified as multiple classes in FR.
-In this section, we will explore the existence of multi-identity samples among the natural face images.
-Table 3: Comparisons of ASR (%) results of dodging attack and impersonation attacks on the LFW dataset. The models in the second row are the victim models.
-ASRd ASRi
-Attack IR152 FaceNet MF
-Baseline 2.2 18.2 99.2 25.6
-Lagrangian 3.9 26.5 100.0 26.0
-Fine-tuning 4.2 26.3 100.0 25.6
-RZ 0.6 4.4 94.8 15.1
-Pruning 3.4 24.8 100.0 25.4
-Adv-Pruning 5.4 32.5 100.0 26.3
-We randomly select negative face pairs from the entire LFW dataset.
-Subsequently, we use MF as our FR model to extract the embeddings of the face images in each face pair and calculate the cosine similarity between the two images.
-If the cosine similarity surpasses the threshold, both images in the pair are classified as belonging to multiple identities, indicating that they are multi-identity samples.
-Our findings demonstrate the presence of multi-identity samples among the benign face images, as illustrated in the bottom left of Fig.
-1.
-The multi-identity samples in Fig.
-1 closely resemble the appearances of the identities they are classified into.
-To analyze the cause of this phenomenon, we need to consider the properties of both the multi-identity samples and the FR model.
-Commonly-used FR models are well-trained and capable of correctly classifying the majority of benign face images.
-However, there are some benign face images that the FR model fails to classify accurately, and these samples are referred to as hard samples [21; 72].
-Multi-identity samples are a specific type of hard sample known as hard negative samples.
-Typically, hard negative samples exhibit a similar appearance [67], indicating that the multi-identity samples among the benign face images share a resemblance.
-Universality of Multi-identity Samples among Adversarial Face Examples: The previous impersonation methods craft adversarial face examples only use the impersonation loss $\mathcal{L}^{i}$.
-In this section, we will investigate the ratio of multi-identity samples and evaluate the effectiveness of previous impersonation methods in terms of dodging attacks.
-We utilize the Multi-identity Sample Ratio (MSR) to gauge the proportion of multi-identity samples in the adversarial face examples capable of executing successful impersonation attacks.
-These multi-identity samples can be recognized as both the attacker and victim identities in the setting of our paper.
-The detailed calculation method of MSR is in the appendices.
-We evaluate the MSR, impersonation ASR and dodging ASR using MF as the surrogate model on the LFW dataset and the results are demonstrated in the appendices.
-The results demonstrates that most of the crafted adversarial face examples are multi-identity samples in the black-box setting.
-This indicates that most adversarial face examples generated through previous impersonation attacks are unable to attain a successful dodging attack in the black-box setting.
-Nevertheless, in the white-box setting, the majority of adversarial face examples capable of executing successful impersonation attacks also demonstrate success in dodging attacks.
-To analyze the reason, we consider the metric used to determine whether two face images belong to the same identity in FR.
-Since FR is an open-set task, we rely on the distance in the embedding space to make decisions.
-The top of Fig.
-1 illustrates two decision boundaries for each FR model, one for attacker identity and one for victim identity.
-In the white-box setting, if we generate adversarial examples using $\mathcal{L}^{i}$, these adversarial examples can penetrate a space where they are recognized as the victim identity rather than the attacker identity.
-However, the decision boundary of the black-box model differs from that of the surrogate model.
-In the black-box setting, most adversarial examples are found between the decision boundaries of the black-box model, resulting in the majority of adversarial examples crafted using $\mathcal{L}^{i}$ being multi-identity samples.
-This demonstrates that the adversarial face examples are positioned near the decision boundary in the black-box setting.
-The Trade-off Between the Impersonation Attacks and Dodging Attacks: Owing to the inherent conflict during the optimization process of impersonation and dodging losses in the black-box setting, there exists a trade-off between impersonation and dodging performance.
+More ablation studies are in the supplementary.
+Figure 6. Performance of dodging ASR across different JPEG Q values. (a) Results using ArcFace as the victim model. (b) Results using CurricularFace as the victim model.
+4.4. Analytical Study
+Owing to the inherent conflict during the optimization process of impersonation and dodging losses in the black-box setting, there exists a trade-off between impersonation and dodging performance.
 We craft adversarial face examples using Lagrangian attack and our proposed Adv-Pruning on LFW based on DI.
-The average black-box results are demonstrated in the left plot of Fig.
-7.
-Figure 7: (a) The trade-off between the ASR (%) of impersonation attack and dodging attack of adversarial examples. (b) The dodging ASR (%) in different sparsity ratios.
+The average black-box results are demonstrated in the left plot of Figure 7.
+Figure 7. (a) The trade-off between the ASR (%) of impersonation attack and dodging attack of adversarial examples. (b) The dodging ASR (%) in different sparsity ratios.
 The results illustrate that our proposed method can reduce the trade-off between impersonation and dodging performance in the black-box setting.
@@
 These operations effectively enhance the dodging attack performance while maintaining the impersonation attack performance, ultimately mitigating the trade-off.
-5 Conclusion
-In this paper, we reveal the universality of multi-identity samples among adversarial face examples crafted by previous impersonation attacks and the success of an impersonation attack may not necessarily imply the success of dodging attacks on FR systems in the black-box setting.
+The analytical study of multi-identity samples among the benign face images and universality of multi-identity samples among adversarial face examples are in the supplementary.
+5. Conclusion
+In this paper, we delve into the issue of multi-identity samples among adversarial face examples.
+Our research reveals the universality of multi-identity samples among adversarial face examples crafted by previous impersonation attacks and the success of an impersonation attack may not necessarily imply the success of dodging attacks on FR systems in the black-box setting.
 In order to improve dodging performance without compromising impersonation performance, we proposed a novel attack, namely Adv-Pruning.
@@
 Extensive experiments demonstrate the effectiveness of our proposed method.
-Appendix A The Pseudo-code of Adversarial Pruning Attack Method
-The pseudo-code of our proposed method based on the Lagrangian attack is illustrated in Algorithm 1.
-Algorithm 1 Adversarial Pruning Attack Based on Lagrangian
-0:  Negative face image pair $\{x^{s},x^{t}\}$, the number of the iterations of Priming stage $n$, the maximum number of iterations $m$, maximum allowable perturbation magnitude $\epsilon$, the surrogate FR model $\mathcal{F}$, the step size $\beta$ and $\gamma$.
-0:  An adversarial face example $x^{adv}_{m}$
-1:  $x^{adv}_{0}=x^{s}$
-2:  for $t=1,...,n$ do
-3:     $\mathcal{L}^{p}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Priming Stage
-4:     $x^{adv}_{t}=x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)$
-5:     $x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{t}\right)$
-6:  end for
-7:  $\mathcal{I}=|x^{adv}_{n}-x^{s}|$triangleright Pruning Stage
-8:  $\mathcal{Q}={\rm Sort}\left(\Psi\left(\mathcal{I}\right)\right)$
-9:  $\mathcal{W}=\mathcal{Q}\left[:\kappa s\right]$
-10:  Get $\mathcal{M}$ by Eq. 8.
-11:  $x^{adv}_{n}=x^{s}+\left(x^{adv}_{n}-x^{s}\right)\odot\mathcal{M}$
-12:  $\mathcal{A}=1-\mathcal{M}$
-13:  for $t=n+1,...,m$ do
-14:     $\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Restoration Stage
-15:     $x^{adv}_{t}=x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)$
-16:     $p=\mathcal{A}\odot\left(x^{adv}_{t}-x^{adv}_{n}\right)$
-17:     $x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{n}+p\right)$
-18:  end for
-Appendix B The Optimization Objective of the Impersonation and Dodging Attacks
-The objective of impersonation attacks in our research is to craft an adversarial example $x^{adv}$ that causes the model $\mathcal{F}^{vct}$ to incorrectly classify it as the target sample $x^{t}$, while simultaneously maintaining a high degree of visual similarity between $x^{adv}$ and the original sample $x^{s}$.
-To be more specific, the objective can be expressed as follows:
-$\begin{gathered}x^{adv}=\mathop{\arg\min}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (13)
-where $\mathcal{D}$ refers to a predefined distance metric, while $\epsilon$ specifies the maximum magnitude of permissible perturbation.
-In contrast, the dodging attacks introduced in this study aim to prevent the model $\mathcal{F}^{vct}$ from correctly recognizing $x^{adv}$ as the source sample $x^{s}$.
-This is achieved while maintaining a visually similar appearance between the adversarial example $x^{adv}$ and the original input $x^{s}$.
-In a manner similar to impersonation attacks, the objective of dodging attacks can be formulated as follows:
-$\begin{gathered}x^{adv}=\mathop{\arg\max}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (14)
-Appendix C The proof of the supremum of the absolute variance
-In this section, we will provide the proof of the supremum of the absolute variance of the output of $\mathcal{F}$ before and after pruning. The absolute variance can be expressed as follows:
-$\rho=\|\mathcal{F}\left(x+\delta\right)-\mathcal{F}\left(x\right)\|$ (15)
-[…]
-Let us revisit an early hypothesis posited by Goodfellow et al.
-[14], suggesting that the linear nature of modern DNNs, resembling linear models trained on the same dataset, is the underlying cause of adversarial examples and their surprising transferability [15].
-Based on the hypothesis, the FR model $\mathcal{F}$ can be represented in the following linear form:
-$\mathcal{F}\left(x\right)=\omega x+b$ (16)
-where $\omega$ is the weight vector, and $b$ is the bias term.
-Hence, the absolute variance of the output of the function $\mathcal{F}$ can be articulated as:
-$\begin{split}\rho&=\|\mathcal{F}\left(x+\delta\right)-\mathcal{F}\left(x\right)\|\\ &=\|\omega\left(x+\delta\right)+b-\left(\omega x+b\right)\|\\ &=\|\omega\delta\|\leq\|\omega\|\|\delta\|=\|\omega\|\mathcal{I}\\ \end{split}$ (17)
-The supremum of the absolute variance is given by $\lVert\omega\rVert\mathcal{I}$.
-Within the context of gradient-based adversarial attacks, the parameters of the neural networks are typically held constant.
-Consequently, the vector $\omega$ can be treated as a constant.
-From this, we can deduce that the supremum of the absolute variance in the output of $\mathcal{F}$ is directly proportional to the priority measure $\mathcal{I}$.
-It follows that adversarial perturbations of greater magnitude are more prioritized, as they exert a more pronounced influence on the output of the face recognition model $\mathcal{F}$.
-Appendix D The Calculation Method of the Attack Success Rate
-Due to the different objectives of the impersonation and dodging attacks, the calculation methods for the ASR also vary between the two attack types.
-When it comes to impersonation attacks on FR, the ASR can be computed as:
-${\rm ASR}^{i}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)<t^{i}\right)}{N_{p}}$ (18)
-where $N_{p}$ refers to the total number of face pairs and $t^{i}$ represents the impersonation attack threshold. For dodging attacks on FR systems, the ASR can be computed via the following formula:
-${\rm ASR}^{d}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)>t^{d}\right)}{N_{p}}$ (19)
-where the value of $t^{d}$ represents the threshold for dodging attacks.
-Appendix E The Calculation Method of the Multi-identity Sample Ratio
-Multi-identity Sample Ratio (MSR) is the proportion of multi-identity samples that can be recognized as both the attacker and victim identities in the adversarial face examples capable of executing successful impersonation attacks.
-The detailed calculation method of MSR be expressed as:
-${\rm MSR}=\frac{|\mathcal{S}_{1}\cap\mathcal{S}_{2}|}{|\mathcal{S}_{2}|}$ (20)
-where $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are the set of the adversarial examples that are identified as the attacker and victim identities, respectively.
-Therefore, the $\mathcal{S}_{2}$ is the set of adversarial face examples that can perform a successful impersonation attack.
-Table 4: Comparisons of ASR results (%) of dodging attack and impersonation attacks on the LFW dataset using the negative cosine similarity loss. The models in the second row are the victim models.
-ASRd ASRi
-IR152 FaceNet MF
-Baseline 2.2 16.1 99.4 25.4
-Baseline + Lagrangian 4.0 26.5 100.0 26.0
-Baseline + ours 8.1 39.7 100.0 26.2
-Appendix F Results on Negative Cosine Similarity Loss
-Apart from the normalized $L_{2}$ distance loss (i.e.
-Eq.
-1 and Eq.
-2), the negative cosine similarity loss is also a commonly used loss function in adversarial attacks on FR systems.
-In this section, we will assess the effectiveness of our proposed method regarding the negative cosine similarity loss.
-For impersonation attacks, the negative cosine similarity loss can be expressed as:
-$\mathcal{L}^{i}_{c}=1-\cos\left(\mathcal{F}\left(x^{adv}\right),\mathcal{F}\left(x^{t}\right)\right)$ (21)
-where $\cos\left(e,e^{\prime}\right)$ is the function to calculate the cosine similarity between $e$ and $e^{\prime}$.
-Similarly, the negative cosine similarity loss for dodging attacks can be formulated as follows:
-$\mathcal{L}^{d}_{c}=1+\cos\left(\mathcal{F}\left(x^{adv}\right),\mathcal{F}\left(x^{s}\right)\right)$ (22)
-To evaluate the effectiveness of our proposed method on the negative cosine similarity loss, we use DI as the Baseline to craft adversarial face examples using the negative cosine similarity loss $\mathcal{L}^{i}_{c}$ and $\mathcal{L}^{d}_{c}$ as $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, respectively.
-The performance is demonstrated in Table 4.
-Table 4 demonstrates that the performance of both impersonation and dodging improves after integrating our proposed attack method into the Baseline attack.
-This further highlights the effectiveness of our method regarding the negative cosine similarity loss.
-Table 5: The results (%) on the LFW dataset using MF as the surrogate model. The models on the first row are the victim models.
-ASRd ASRi
-IR152 FaceNet MF
-SA 99.4 45.9 2.1 44.5
-SA + Lagrangian 100.0 49.2 3.4 44.8
-SA + ours 100.0 54.9 4.7 46.4
-Appendix G Results on Sibling-Attack
-Sibling-Attack [29] (abbreviated as SA) is a cutting-edge adversarial attack on FR that demonstrates considerable effectiveness in enhancing the transferability of adversarial attacks on FR.
-SA employs a surrogate model with a two-branch multi-task neural network, which differs from the surrogate models discussed in Section 4.2.
-Consequently, we exclusively consider SA for comparison with our method.
-We aligned the attack setting with the official SA implementation and integrated our proposed Adv-Pruning with SA.
-The FFHQ dataset, provided by the official SA page [^1] https://github.com/Tencent/TFace/tree/Adv-Attack-Defense/security/tasks/Adv-Attack-Defense/Sibling-Attack, is chosen as the dataset for generating adversarial face examples.
-The results are depicted in Table 5.
-Table 5 demonstrates that the incorporation of our proposed method results in enhancements in both dodging and impersonation performance.
-This further demonstrates the effectiveness of our proposed attack method in enhancing the dodging attack performance without compromising the impersonation attack performance.
-Table 6: The results (%) on the LFW dataset using MF as the surrogate model. The models on the first row are the victim models.
-Attack Metric IR152 FaceNet MF
-DI MSR 95.6 75.7 0.8
-ASRi 18.4 32.9 100.0
-ASRd 2.2 18.2 99.2
-VMI MSR 97.0 88.6 0.3
-ASRi 13.6 20.2 100.0
-ASRd 1.0 8.4 99.7
-SSA MSR 98.6 92.8 1.5
-ASRi 13.8 19.6 100.0
-ASRd 0.7 5.9 98.5
-DFANet MSR 98.6 96.6 0.4
-ASRi 7.0 11.9 100.0
-ASRd 0.2 1.5 99.6
-SIA MSR 97.4 83.9 1.6
-ASRi 15.7 26.7 100.0
-ASRd 1.0 10.6 98.4
-BSR MSR 98.1 96.8 15.1
-ASRi 5.4 9.5 100.0
-ASRd 0.4 3.7 84.9
-BPFA MSR 96.2 92.1 2.3
-ASRi 15.8 17.8 100
-ASRd 0.9 4.6 97.7
-Appendix H Results on MSR of Previous Impersonation Attacks
-Following the same experimental setting of Section 4.4, we conduct experiments on various previous impersonation attacks. The results are shown in Table 6.
-As indicated by Table 6, the majority of the crafted adversarial face examples are multi-identity samples in the black-box setting.
-This observation suggests that the majority of adversarial face examples, which are generated through the previous impersonation attacks, fail to achieve a successful dodging attack in the black-box setting.
+Acknowledgements.
+This work was supported in part by the Natural Science Foundation of China under Grant 61972169,62372203 and 62302186, in part by the National Key Research and Development Program of China (2022YFB2601802), in part by the Major Scientific and Technological Project of Hubei Province (2022BAA046, 2022BAA042), in part by the Knowledge Innovation Program of Wuhan-Basic Research, in part by China Postdoctoral Science Foundation 2022M711251.